### v2

# The Holomorphic Extension Property for $k$-du Bois Singularities

Abstract.
Let $X$ be a normal complex algebraic variety. We show that the holomorphic extension property holds in degree $p<\mathrm{codim}_{X}(X_{\mathrm{sing}})$ when $X$ has du Bois singularities, giving an improvement on Flenner’s criterion for arbitrary singularities. As an application, we study the $k$-du Bois definition from the perspective of holomorphic extension and compare how different restrictions on $\mathscr{H}^{0}(\underline{\Omega}_{X}^{p})$ affect the singularities of $X$, where $\underline{\Omega}_{X}^{p}$ is the $p^{th}$-graded piece of the du Bois complex.
1. Introduction
1.1. Holomorphic and Logarithmic Extension
Let $X$ be a normal complex variety with regular locus $U$ with inclusion morphism $j:U\hookrightarrow X$. We say that the holomorphic extension property holds in degree $p$ if the natural inclusion
(1.1) $\pi_{*}\Omega_{\widetilde{X}}^{p}\hookrightarrow\Omega_{X}^{[p]}:=j_{*}\Omega_{U}^{p}$
is an isomorphism for some — and therefore any — resolution of singularities $\pi:\widetilde{X}\to X$.
The holomorphic extension property has been extensively studied for the classes of singularities arising in the minimal model program. For klt and rational singularities, the holomorphic extension property holds for every $p$, see [GKKP11, Theorem 1.4] and [KS21, Corollary 1.8]. Rational singularities consequently satisfy many important properties:
  • Functorial pullback for the sheaves of reflexive differentials $\Omega_{X}^{[p]}$ [Keb13, Theorem 1.3], [KS21, Theorem 1.11].
  • The sheaf $\Omega_{X,h}^{p}$ of $h$-differential $p$-forms agrees with $\Omega_{X}^{[p]}$ [HJ14, Thm. 1], [KS21, Corollary 1.12].
  • The Zariski-Lipman conjecture holds for rational singularities: given a normal complex analytic variety with rational singularities and locally free tangent sheaf, then $X$ is smooth [GKKP11, §6], [KS21, Theorem 1.14].
Holomorphic extension is a weak condition for small $p$: if $\Sigma$ is the singular locus of $X$, then (1.1) is an isomorphism for every $10\leq p<\mathrm{codim}_{X}(\Sigma)-$ 1 by Flenner’s criterion [Fle88, Theorem]. Moreover, if holomorphic extension holds in degree $p$, then holomorphic extension holds in degree $k$ for every $0\leq k\leq p$ [KS21, Theorem 1.4]. For $p\geq\mathrm{codim}_{X}(\Sigma)-1$, extension is less topological and (1.1) can be strict: by definition, strictly log-canonical singularities fail holomorphic extension in degree $\dim X$. Instead, it is better to consider a variant of the inclusion (1.1) which allows for logarithmic poles. We say that a normal variety $X$ satisfies the logarithmic extension property in degree $p$ if the inclusion
(1.2) $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\hookrightarrow\Omega_{X}^{[p]}$
is an isomorphism for some — and therefore any — log-resolution of singularities $\pi:\widetilde{X}\to X$ with log-exceptional divisor $E=\pi^{-1}(\Sigma)$. Log-canonical singularities satisfy the logarithmic extension property for all $0\leq p\leq\dim X$ [GKKP11, Theorem 1.5]. In general, logarithmic extension in degree $p$ implies logarithmic extension in degree $k$ for every $0\leq k\leq p$ [KS21, Theorem 1.5].
1.2. Extension for du Bois Singularities
Let $X$ be a complex algebraic variety. We say that $X$ has du Bois singularities if the natural morphism
(1.3) $\mathscr{O}_{X}\to\underline{\Omega}_{X}^{0}$
is a quasi-isomorphism, where $\underline{\Omega}_{X}^{0}:=\mathrm{gr}_{F}^{0}\underline{\Omega}_{X}^{\bullet}$ is the $0^{th}$-graded piece of the du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$, an object in the category of filtered complexes of $\mathscr{O}_{X}$-modules generalizing the holomorphic de Rham complex of smooth algebraic varieties.
du Bois singularities are an important in the Hodge theory of singularities, as they arise in the construction of Deligne’s mixed Hodge structure [Del74] on the cohomology of algebraic varieties. They also play a fundamental role in the minimal model program, as log-canonical and rational singularities are du Bois, see [KK10, Theorem 1.4] [Kov99, Corollary 2.6], [Sai00, 5.4. Theorem]. We consider then the following questions:
  1. (i)
    If $X$ has du Bois singularities, for which $p$ is the inclusion morphism (1.1) an isomorphism?
  2. (ii)
    If $X$ has du Bois singularities, is the inclusion morphism (1.2) an isomorphism for all $p$?
If $X$ is du Bois and Cohen-Macaulay, the answer to Question (ii) is already known: if $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, then the sheaf $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive by [KSS10, Theorem 1.1] [^1] In fact, a Cohen-Macaulay variety is du Bois if and only if $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive, giving a natural generalization of Kempf’s criterion for rational singularities., and so $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is reflexive for all $0\leq p\leq\dim X$ [KS21, Theorem 1.5]. We can extend the ideas of [KSS10] to give an analytic proof of logarithmic extension for normal du Bois singularities, giving a positive answer to Question (ii):
Theorem 1.1.
If $X$ is a normal complex algebraic variety with at worst du Bois singularities, then logarithmic extension holds in all degrees $0\leq p\leq\dim X$: for any logarithmic resolution of singularities $\pi:\widetilde{X}\to X$, the inclusion $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\hookrightarrow\Omega_{X}^{[p]}$ is an isomorphism.
Theorem 1.1 is key to understanding for which $p$ holomorphic extension holds. Our main result is an extension of Flenner’s criterion to du Bois singularities and an optimal answer to Question (i).
Theorem 1.2.
If $X$ is a normal complex algebraic variety with at worst du Bois singularities and singular locus $\Sigma$, then holomorphic extension holds in degree $0\leq p<\mathrm{codim}_{X}(\Sigma)$: if $\pi:\widetilde{X}\to X$ is any resolution of singularities, the inclusion morphism $\pi_{*}\Omega_{\widetilde{X}}^{p}\hookrightarrow\Omega_{X}^{[p]}$ is an isomorphism.
1.3. Proof of Theorem 1.2 for Isolated Singularities
We note that Theorem 1.2 (and its dependence on Theorem 1.1) has been observed in the literature in special cases. For one, Graf-Kovács observe the Zariski-Lipman conjecture holds for du Bois singularities by demonstrating that the natural inclusion
(1.4) $\pi_{*}\Omega_{\widetilde{X}}^{p}\hookrightarrow\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$
is an isomorphism for $p=1$, where $\pi:\widetilde{X}\to X$ is a log-resolution of singularities, and showing $\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)$ is reflexive [GK14]. By Flenner’s criterion, this is novel exactly when $\mathrm{codim}_{X}(\Sigma)=2$. Theorem 1.2 can also be seen to hold as a corollary of Theorem 1.1 for isolated singularities by an old result of van Straten-Steenbrink: if $j:U\hookrightarrow X$ is the inclusion of the regular locus $U$, there is a differentiation morphism
$d:j_{*}\Omega_{U}^{\dim X-1}/\pi_{*}\Omega_{\widetilde{X}}^{\dim X-1}\to j_{*}\omega_{U}/\pi_{*}\omega_{\widetilde{X}}(E)$
for any log-resolution of singularities $\pi:\widetilde{X}\to X$, which is injective if $X$ has isolated singularities [vSS85, Cor. 1.4].
An obvious idea is to reduce Theorem 1.2 to the case of isolated singularities by cutting $X$ down by successive hyperplane sections. More specifically, let $\pi:\widetilde{X}\to X$ be a log-resolution of singularities with exceptional divisor $E$. If $H$ is a very general hyperplane section of $X$, then there is an induced log-resolution of singularities $\pi|_{H}:\widetilde{H}\to H$, and we let $E|_{H}$ be the induced exceptional divisor. For $p\geq 1$, consider the commutative diagram
[…]
Recalling that general hyperplane sections preserve the du Bois property, an inductive hypothesis would imply that the left and right vertical morphisms are isomorphisms — this is exactly the approach used in [GK14] for $p=1$, but they need to use the negativity lemma [GKK10, Proposition 7.5], which seems particularly special for $p=1$. Induction seems insufficient in proving $\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is an isomorphism, and the reflexivity of the sheaf $(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p}(\log E|_{H})$ is not sufficient to prove the reflexivity of $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$. We require additional input.
1.4. Extension Criterion and Hodge Modules
To prove Theorem 1.2, we consider the following well-known interpretation of extension for coherent sheaves:
Proposition 1.3.
[KS21, Corollary 6.2] Let $Y$ be a complex manifold and $\mathscr{F}$ a coherent sheaf of $\mathscr{O}_{Y}$-modules. If $\mathrm{Supp}~{}\mathscr{F}$ has pure dimension $n$, the following are equivalent:
  1. (i)
    Section of $\mathscr{F}$ extend uniquely across any subset $A\subset Y$ with $\dim A\leq n-2$.
  2. (ii)
    For every $k\geq-n+1$ $\dim\mathrm{Supp}~{}R^{k}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathscr{F},\omega_{Y}^{\bullet})\leq-(k+2$), where $\omega_{Y}^{\bullet}$ is the dualizing complex of $Y$.
We will often apply this to a singular variety $X$ of dimension $n$ by considering a local embedding $X\subset Y$ into a smooth complex manifold $Y$.
To highlight the idea, let $X$ be a normal variety of dimension $n$ with at worst isolated du Bois singularities, which implies that the cohomology sheaves $R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$ for $i>0$ for any log-resolution of singularities $\pi:\widetilde{X}\to X$. By Grauert-Rimenschneider vanishing, there is an induced quasi-isomorphism
(1.6) $\pi_{*}\omega_{\widetilde{X}}(E)\cong\mathbf{R}\pi_{*}\omega_{\widetilde{X}}(E),$
and Grothendieck duality immediately implies that
$R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(\pi_{*}\omega_{\widetilde{X}}(E),\omega_{X}^{\bullet})[-n]\cong R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$
for $j>0$. By Proposition 1.3, $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive. We emphasize that this vanishing is much stronger than logarithmic extension and is only equivalent if $X$ is Cohen-Macaulay. Theorem 1.1 follows more generally from the following lemma:
Lemma 1.4.
If $X$ is a normal complex algebraic variety with a log-resolution of singularities $\pi:\widetilde{X}\to X$, then $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{0})\cong\pi_{*}\omega_{\widetilde{X}}(E)$, where $\mathbb{D}_{X}$ is the Grothendieck duality functor.
For holomorphic extension, the problem is more subtle. We need to extend Proposition 1.3 to the level of complexes:
Proposition 1.5.
[KS21, Proposition 6.4] Let $Y$ be a complex manifold, let $A\subset Y$ be a complex subspace, and let $K\in\mathrm{D}_{\mathrm{coh}}^{b}(\mathscr{O}_{Y})$ be a complex with $\mathscr{H}^{j}K=0$ for $j<0$. If
$\dim(A\cap\mathrm{Supp}~{}R^{k}\mathscr{H}om_{\mathscr{O}_{Y}}(K,\omega_{Y}^{\bullet}))\leq-(k+2)$
for every $k\in\mathbb{Z}$, then the sections of $\mathscr{H}^{0}K$ extend uniquely across $A$.
The naive approach is to consider the support of the sheaves $R^{j}\pi_{*}\Omega_{\widetilde{X}}^{\dim X-p}$ for $p=\mathrm{codim}_{X}(\Sigma)$, but this seems hopeless: for instance, $R^{1}\pi_{*}\Omega_{\widetilde{X}}^{1}$ is non-zero for ADE surface singularities — these singularities satisfy holomorphic extension in all degrees, either classically or by [KS21, Corollary 1.8]. Instead, we use the brilliant approach of Kebekus-Schnell and the theory of Hodge modules.
Let $X$ be a normal complex algebraic variety with du Bois singularities. There are objects $K_{p},K_{p}^{\prime}\in\mathrm{D}_{\mathrm{coh}}^{b}(\mathscr{O}_{X})$ such that $K_{p}$ defines a sub-object of $\mathbf{R}\pi_{*}\Omega_{\widetilde{X}}^{p}$ for every $p$, and
$\pi_{*}\Omega_{\widetilde{X}}^{p}\cong\mathscr{H}^{0}K_{p},\quad\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\cong\mathscr{H}^{0}K_{p}^{\prime}.$
The objects $K_{p}$ are defined by the intersection cohomology complex and its data as a pure Hodge module; the objects $K_{p}^{\prime}$ are determined by the data of a mixed Hodge module, which we refer to as the logarithmic mixed Hodge module. Both are uniquely determined as extensions of the trivial Hodge module $\mathbb{Q}_{\mathrm{reg}}$.
Assuming that $X$ has du Bois singularities, Theorem 1.1 and [KS21, §9] imply a family of support conditions
$\dim\mathrm{Supp}~{}R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(K_{p}^{\prime},\omega_{X}^{\bullet})\leq-(k+2).$
The key is to relate this support condition to $K_{p}$. In particular, an inductive argument on $\dim\Sigma$ allows us to prove the following, which gives a stronger result than Theorem 1.2:
Theorem 1.6.
Let $X$ be a normal complex algebraic variety with at worst du Bois singularities. For $p=\mathrm{codim}_{X}(\Sigma)-1$, we have
$\dim\mathrm{Supp}~{}\mathscr{H}^{j+n-p}K_{\dim X-p}\leq-(j+2),$
where
$K_{p}:=\mathrm{gr}_{-p}^{F}\mathrm{DR}(IC_{X})[p-n]$
is the $p^{th}$-graded piece of the intersection cohomology Hodge module with its induced Hodge filtration.
1.5. Log Forms to Holomorphic Forms
It is known that Flenner’s criterion for holomorphic extension is optimal for general singularities; in particular, there is a variety $X$ with non-du Bois singularities for which holomorphic extension fails in degree $p=\mathrm{codim}_{X}(\Sigma)-1$. We consider then a weakening of holomorphic extension via 1.4: if a differential form on $X_{\mathrm{reg}}$ extend with at worst log-poles, does it already extend holomorphically? Of course, (1.4) is surjective whenever holomorphic extension holds in degree $p$, and so this holds for any variety with $p<\mathrm{codim}_{X}(\Sigma)-1$. We consider again what happens for $p=\mathrm{codim}_{X}(\Sigma)-1$. This generalizes the discussion in §5.2.
Theorem 1.7.
Let $X$ be a normal variety with singular locus $\Sigma$. The inclusion morphism (1.4) is an isomorphism for $p=\mathrm{codim}_{X}(\Sigma)-1$.
1.6. Holomorphic Extension for $k$-du Bois Singularities
du Bois singularities are of fundamental interest to algebraic geometers as a general class of singularities for which deformation theory and Hodge theory can be studied. It is interesting to consider $k$-du Bois singularites, the class of singularities which restricts the higher graded pieces of the du Bois complex
We say that a normal complex algebraic variety $X$ has weakly (or pre-) $k$-du Bois singularities if the cohomology sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}=0$ for $j>0$ and $0\leq p\leq k$, where $\underline{\Omega}_{X}^{p}:=\mathrm{gr}_{F}^{p}\underline{\Omega}_{X}^{\bullet}[p]$. Alternatively, $X$ has weakly $k$-du Bois singularities if the natural map
$\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$
is a quasi-isomorphism. This notion has been studied in [SVV23] and is a generalization of the notion of $k$-du Bois singularities appearing in [MOPW21], [JKSY21], and [MP22].
An interesting question is then to what extent Theorem 1.2 can be improved for weakly $k$-du Bois singularities. For hypersurfaces, the answer is known: a variety $X$ with hypersurface singularities satisfies holomorphic extension in degree $\dim X$ if and only if $X$ is 1-du Bois, which means that $X$ is du Bois and the natural map
$\Omega_{X}^{1}\xrightarrow{\sim}\underline{\Omega}_{X}^{1}$
is a quasi-isomorphism, where $\Omega_{X}^{1}$ is the sheaf of Kähler 1-forms. This seems to be rather special for hypersurface (or more generally complete intersection) singularities, as we can write down a Cohen-Macaulay variety with weakly $k$-du Bois singularities and $k>0$ in any dimension for which Theorem 1.2 is still optimal (see Example 6.1). The algebraic properties of the sheaf $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ appear to be independent of the vanishing — or non-vanishing — of the sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}$. Instead, we consider what happens when $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is already reflexive. One result in this direction is the following:
Theorem 1.8.
Let $X$ be a normal variety with weakly $k$-du Bois singularities, and let $\pi:\widetilde{X}\to X$ be a resolution of singularities. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then $\pi_{*}\Omega_{\widetilde{X}}^{k+1}$ is reflexive.
1.7. Higher Rational Singularities and Reflexivity of $\Omega_{X}^{p}$
Let $X$ be a normal complex algebraic variety. Following [MOPW21] and [JKSY21], we say that $X$ has $k$-du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $p$, where $\Omega_{X}^{p}$ is the sheaf of Kähler $p$-forms. As a generalization of du Bois singularities, this is very natural, as it comes from a morphism of complexes $\Omega_{X}^{\bullet}\to\underline{\Omega}_{X}^{\bullet}$. However, this definition is rather restrictive: in the case of hypersurface singularities, this forces $\Omega_{X}^{p}$ to be reflexive for every $0\leq p\leq k$ and restricts the codimension of the singularities.
Instead, one can consider what happens when we restrict the duals of the higher du Bois complexes. We say that has weakly $k$-rational singularities if
$\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$
for each $0\leq p\leq k$, where $\mathbb{D}_{X}$ is the Grothendieck duality functor. Note that there is a quasi-isomorphism $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n})\cong\mathbf{R}\pi_{*}\mathscr{O}_{\widetilde{X}}$, whence 00-rational is the same as rational singularities for normal complex varieties. This is a generalization of the $k$-rational definition established for complete intersections, which requires the natural morphism $\Omega_{X}^{p}\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ to be an isomorphism for $0\leq p\leq k$.
As was observed in [MOPW21], the $k$-du Bois definition restricts the codimension of the singularities of a hypersurface for large $k$, and $X$ is smooth if and only if $k>\frac{1}{2}\dim X$. We generalize this result for $k$-rational singularities. We remark that this appears to be the correct generalization, as $k$-du Bois singularities are $(k-1)$-rational for hypersurfaces by [FL22b].
Theorem 1.9.
If $X$ is a normal complex algebraic variety with $k$-rational and Gorenstein singularities, then $k\leq\frac{1}{2}\mathrm{codim}_{X}(\Sigma)$.
1.8. Hodge Theory of Weakly $k$-Rational Singularities
As we mentioned above, the du Bois complex is a Hodge-theoretic object arising from Deligne’s mixed Hodge theory. If $X$ is a proper variety, then the Hodge filtration is the one induced by the $E_{1}$-spectral sequence
$E_{1}^{p,q}=\mathbb{H}^{q}(X,\underline{\Omega}_{X}^{p})\Rightarrow H^{p+q}(X,\mathbb{C}).$
If $X$ is weakly $k$-du Bois, we get a decomposition
$H^{k}(X,\mathbb{C})=\bigoplus_{p+q=k}H^{q}(X,\mathscr{H}^{0}\underline{\Omega}_{X}^{p}).$
In general, restricting the du Bois complex does not affect the weight filtration on the cohomology, as $\underline{\Omega}_{X}^{\bullet}$ does not usually admit the structure of a Hodge module. To get pure Hodge modules, we consider again $k$-rational definition.
Theorem 1.10.
If $X$ is a normal complex algebraic variety with weakly $k$-rational singularities, then the Hodge filtration on $H^{k}(X,\mathbb{Q})$ induces a pure Hodge structure.
1.9. Acknowledgements
Part of this paper is contained in the author’s PhD thesis. I want to thank Benjamin Bakker and Nicolas Addington for conversations regarding this topic. I want to thank Mihnea Popa, Wanchun Shen, and Duc Vo for comments on an early draft of this paper. Finally, I thank Sung Gi Park, who independently proved Theorem 1.2 and Theorem 1.1 in [Par23], for conversations on this topic.
2. The du Bois Complex
2.1. Notation
Throughout, we let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$. We will use $\pi:\tilde{X}\to X$ to denote a strong and projective resolution of singularities.
If $\pi:\tilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, we let $\Omega_{\tilde{X}}^{p}(\log E)$ be the sheaf of log $p$-forms, and we let
$\Omega_{\tilde{X}}^{p}(\log E)(-E):=\Omega_{\tilde{X}}^{p}(\log E)\otimes\mathscr{I}_{E}.$
For any local embedding $X|_{V}\hookrightarrow Y$ into a smooth complex variety, we define the sheaf of Kähler differentials
$\Omega_{X}^{1}|_{V}:=\Omega_{Y}^{1}/\langle df_{1},df_{2},...,df_{m}\rangle,$
where $\Omega_{Y}^{1}$ is the sheaf of holomorphic 1-forms on $Y$ and $f_{1},...,f_{m}$ are some defining equations for the open set $V$. We let $\Omega_{X}^{p}:=\wedge^{p}\Omega_{X}^{1}$.
If $f:X^{\prime}\to X$ is a morphism, we write $\mathbf{R}f_{*}$ for the derived pushforward, and $\mathbf{R}\mathscr{H}om$ for the derived hom. We will use $H^{k}$ for the cohomology of a sheaf, $\mathscr{H}^{k}$ for the cohomology sheaf of a complex, and $\mathbb{H}^{k}$ for hypercohomology of a complex. Finally, we let $\mathbb{D}_{X}(-):=\mathbf{R}\mathscr{H}om_{\mathscr{O}_{X}}(-,\omega_{X}^{\bullet})$ be the Grothendieck duality functor.
2.2. The du Bois Complex
Let $X$ be a complex algebraic variety. The du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$ is an object in the derived category of filtered complexes generalizing the holomorphic de Rham complex for algebraic varieties over $\mathbb{C}$. We denote its graded pieces by $\underline{\Omega}_{X}^{p}:=\mathrm{gr}_{F}^{p}\underline{\Omega}_{X}^{\bullet}[p]$. Studied by du Bois [DB81] from Deligne’s construction of the mixed Hodge structure [Del74], the du Bois complex is constructed by simplicial or cubical hyperresolutions of $X$. We will not need this construction in this paper but will only use its formal consequences. The interested reader can consult [PS08] for a good treatment of this construction.
Theorem 2.1.
For $X$ a complex scheme of finite type and $\underline{\Omega}_{X}^{\bullet}$ its du Bois complex, we have
  1. (i)
    [DB81, §3.2] $\underline{\Omega}_{X}^{\bullet}\cong_{\mathrm{qis}}\mathbb{C}_{X}$.
  2. (ii)
    [DB81, (3.2.1)]If $f:Y\to X$ is a proper morphism of finite type schemes, then there is a morphism $f^{*}:\underline{\Omega}_{X}^{\bullet}\to\mathbf{R}f_{*}\underline{\Omega}_{Y}^{\bullet}$ in $\mathscr{D}_{\mathrm{filt}}^{b}(X)$.
  3. (iii)
    [DB81, 3.10 Corollaire] If $U\subset X$ is an open subscheme then $\underline{\Omega}_{X}^{\bullet}|_{U}\cong_{\mathrm{qis}}\underline{\Omega}_{U}^{\bullet}$.
  4. (iv)
    [GAPGP06, V, 3.6] $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\underline{\Omega}_{X}^{p}\leq\dim X-j$ for $0\leq j\leq\dim X$.
  5. (v)
    [DB81, §3.2] There is a natural morphism $\Omega_{X}^{\bullet}\to\underline{\Omega}_{X}^{\bullet}$, where $\Omega_{X}^{\bullet}$ is the complex of Kähler differentials. Moreover, this morphism is a quasi-isomorphism if $X$ is smooth.
  6. (vi)
    [Kov11, Proposition 2.6], [SVV23, Lemma 3.2] If $H\subset X$ is a general member of a basepoint free linear system, then $\underline{\Omega}_{H}^{\bullet}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{\bullet}\otimes\mathscr{O}_{H}$.
  7. (vii)
    [DB81, 4.11 Proposition] There is an exact triangle
    $\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{\Sigma}^{p}\oplus\mathbf{R}\pi_{*}\Omega_{\tilde{X}}^{p}\to\mathbf{R}\pi_{*}\underline{\Omega}_{E}^{p}\xrightarrow{+1}$
    where $\pi:\tilde{X}\to X$ is a morphism of complex schemes of finite type with singular locus $\Sigma\subset X$ and $E=\pi^{-1}(\Sigma)$. In particular, $\underline{\Omega}_{X}^{n}\cong_{\mathrm{qis}}\mathbf{R}\pi_{*}\omega_{\tilde{X}}\cong_{\mathrm{qis}}\pi_{*}\omega_{\tilde{X}}$ if $\pi:\tilde{X}\to X$ is a resolution of singularities.
  8. (viii)
    [DB81, 4.5 Théorème] If $X$ is a proper variety, there is a spectral sequence
    $E_{1}^{p,q}:=\mathbb{H}^{q}(X,\underline{\Omega}_{X}^{p})\Rightarrow H^{p+q}(X,\mathbb{C})$
    which degenerates at $E_{1}$ for every $p,q$. Moreover, the filtration induced by this degeneration is equal to the Hodge filtration on the underlying mixed Hodge structure.
  9. (ix)
    [Kov11, §3.C] If $\pi:\tilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, there is a right triangle
    $\mathbf{R}\pi_{*}\Omega_{\tilde{X}}^{p}(\log E)(-E)\to\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{\Sigma}^{p}\xrightarrow{+1}$
    which is independent of the choice of $\pi$.
2.3. du Bois, $k$-du Bois, and $k$-rational Singularities
Let $X$ be a complex algebraic variety. We review classes of singularities which are associated to the complexes $\underline{\Omega}_{X}^{p}$.
Definition 2.2.
Let $X$ be a complex algebraic variety.
  1. (i)
    We say that $X$ has du Bois singularities if the natural map $\mathscr{O}_{X}\to\underline{\Omega}_{X}^{0}$ is a quasi-isomorphism.
  2. (ii)
    We say that $X$ has $k$-du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq k$.
  3. (iii)
    We say that $X$ has $k$-rational singularities if the natural map $\Omega_{X}^{p}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ is a quasi-isomorphism for each $0\leq p\leq k$.
  4. (iv)
    We say that $X$ has weakly $k$-du Bois singularities, or pre $k$-du Bois singularities, if the natural map $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq k$.
  5. (v)
    We say that $X$ has weakly $k$-rational singularities, or pre $k$-rational singularities, if the natural map $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ for each $0\leq p\leq k$.
Here are some general properties of (weakly) $k$-du Bois (resp. (weakly) $k$-rational singularities).
  • By Theorem 2.1(vi), if $H$ is a general member of a basepoint linear system of a variety $X$ with (weakly) $k$-du Bois singularities, then $H$ also has (weakly) $k$-du Bois singularities (resp. (weakly) $k$-rational singularities).
  • Rational and log-canonical singularities are du Bois, see [Kov99, Corollary 2.6], [Sai00, 5.4. Theorem], and [KK10, Theorem 1.4].
  • If $X$ is a normal variety and $f:Y\to X$ is a finite dominant map from a variety $Y$ with rational and weakly $k$-du Bois singularities, then $X$ also has rational [Kov00, Theorem 1] and weakly $k$-du Bois (resp. weakly $k$-rational) singularities [SVV23, Proposition 4.2].
  • If $X$ has simple normal crossing singularities, then $X$ is weakly $k$-du Bois since
    $\underline{\Omega}_{X}^{p}\cong\Omega_{X}^{p}/\mathrm{tor}$
    for each $0\leq p\leq n$.
  • If $X$ has rational singularities, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\cong\Omega_{X}^{[p]}$ for each $p$, which follows from [KS21] and [HJ14]. In particular, a complex algebraic variety $X$ with rational singularities is weakly $k$-du Bois (resp. $k$-rational) if and only if $\Omega_{X}^{[p]}\to\underline{\Omega}_{X}^{p}$ (resp. $\Omega_{X}^{[p]}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$) is a quasi-isomorphism for each $0\leq p\leq k$.
There is a nice characterization due to Schwede which wish to mention. For $X$ a reduced and separated complex scheme of finite type, there exists a (local) embedding $\iota:X\to Y$ of $X$ into a smooth scheme $Y$. Let $\pi:\widetilde{Y}\to Y$ be an embedded resolution of $X$ which is an isomorphism outside of $X$, and let $\overline{X}=\pi^{-1}(X)_{\mathrm{red}}$ be the reduced preimage.
Proposition 2.3.
[Sch07, Theorem 4.6] A complex algebraic variety $X$ has du Bois singularities if and only if the natural map $\mathscr{O}_{X}\to\mathbf{R}\pi_{*}\mathscr{O}_{\overline{X}}$ is a quasi-isomorphism.
As we mentioned, the du Bois complex depends on the existence of hyperresolutions of singularities. For lack of a reference, we remark that the conditions of Schwede’s criterion hold in the analytic category, as embedded resolutions of singularities exist by [BM97].
Definition 2.4.
Let $X$ be a complex analytic variety. We say that $X$ has du Bois singularities if for (local) embedding $X\subset Y$ and any embedded resolution of singularities $\pi:\widetilde{Y}\to Y$ which is an isomorphism outside of $X$, the canonical morphism
$\mathscr{O}_{X}\to\mathbf{R}\pi_{*}\mathscr{O}_{\overline{X}}$
is a quasi-isomorphism, where $\overline{X}:=\pi^{-1}(X)$ is the reduced preimage.
Unfortunately, Schwede’s criterion fails for the higher graded pieces, as the proof depends on the fact that simple normal crossing singularities are (0-) du Bois. We note that a hyperresolution-free description of the complexes $\underline{\Omega}_{X}^{p}$ has been given in [Ham23].
3. Hodge Modules and Differentials on the Resolution
3.1. Mixed Hodge Modules
(Mixed) Hodge modules are generalizations of variations of (mixed) Hodge structures in the presence of singularities. We review some definitions concerning mixed Hodge modules following [KS21]. The interested reader may also refer to [Sch19] for more details.
3.1.1. Pure and Mixed Hodge Modules
Let $Y$ be a smooth complex manifold of dimension $d$ (for example, let $Y$ be a local embedding of a complex algebraic variety $X$ of codimension $c=d-n$). A pure Hodge module on $Y$ is an object $M=(\mathcal{M},F_{\bullet},\mathrm{rat}~{}M)$ consisting of:
  1. (i)
    A regular holonomic left $\mathscr{D}_{Y}$-module $\mathcal{M}$, where $\mathscr{D}_{Y}$ is the sheaf of differential operators on $Y$;
  2. (ii)
    An increasing good filtration $F_{\bullet}\mathcal{M}$ of coherent $\mathscr{O}_{Y}$-modules, called the Hodge filtration, which is compatible with the $\mathscr{D}_{Y}$-module structure:
    $F_{p}\mathcal{M}\cdot F_{q}\mathscr{D}_{Y}\subset F_{p+q}\mathcal{M},$
    and $\mathrm{gr}_{F}^{\bullet}$ is coherent $\mathrm{gr}_{F}^{\bullet}\mathscr{D}_{Y}$.
  3. (iii)
    A perverse sheaf $\mathrm{rat}~{}M$ of $\mathbb{Q}$-vector spaces satisfying
    $\mathrm{rat}~{}M\otimes\mathbb{C}\cong\mathrm{DR}(\mathcal{M}),$
    where $\mathrm{DR}$ is the de Rham complex
    $\mathrm{DR}(\mathcal{M}):=\big{[}\mathcal{M}\to\Omega_{Y}^{1}\otimes_{\mathscr{O}_{Y}}\mathcal{M}\to...\to\Omega_{Y}^{d}\otimes_{\mathscr{O}_{Y}}\mathcal{M}\big{]}[d]$
    associated to the $\mathscr{D}_{Y}$-module $\mathcal{M}$. In particular, we have the support condition
    (3.1) $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{DR}(\mathcal{M})\leq-j.$
Example 3.1.
If $Y$ is a smooth complex manifold of dimension $d$, the locally constant sheaf $\mathbb{Q}_{Y}$ admits the structure of a variation of pure Hodge structures. It therefore inherits a Hodge module structure. The underlying left $\mathscr{D}_{Y}$-module is simply $\mathscr{O}_{Y}$, considered as a subsheaf of $\mathscr{E}nd_{\mathbb{C}_{Y}}(\mathscr{O}_{Y})$. The Hodge filtration is the trivial filtration $F_{p}\mathscr{O}_{Y}=\mathscr{O}_{Y}$ for $p\geq 0$ (and zero otherwise). The perverse structure gives the identification
[…]
One can generalize the previous construction to any variation of pure Hodge structure. Conversely, pure Hodge modules satisfy decomposition by strict support [Sai88, §5]: a Hodge module is completely determined by a collection of variations of pure Hodge structures $M_{Y_{i}}$ supported on a stratification $\left\{Y_{i}\right\}$ of $Y$, of weights $w-\dim Y_{i}$, for some $w$. We refer $w$ as the weight of the Hodge module. Because of this, we can define the category $\mathrm{HM}(Y,w)$ of polarized Hodge modules of weight $w$ by inducing a polarization from the data $M_{Y_{i}}$ of polarized variations of pure Hodge structures.
A mixed Hodge module on $Y$ is an object $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)$ consisting of a $\mathscr{D}_{Y}$-module structure $\mathcal{M}$, a perverse structure $\mathrm{rat}~{}M$, a decreasing good filtration $F_{\bullet}$ on $\mathcal{M}$, and an increasing weight filtration $W_{\bullet}$ on these structures such that the graded pieces
$\mathrm{gr}_{k}^{W}M=(\mathrm{gr}_{k}^{W}\mathcal{M},\mathrm{rat}~{}\mathrm{gr}_{k}^{W}M,F_{\bullet})$
are pure Hodge modules. We call $M$ a graded-polarizable mixed Hodge module if further $\mathrm{gr}_{k}^{W}M$ is polarizable for each $k$. We refer to the category of graded-polarizable mixed Hodge modules as $\mathrm{MHM}(Y)$.
3.1.2. The Dual Mixed Hodge Module
Let $Y$ be a complex manifold of dimension $d$, and fix a mixed Hodge module $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$.
We define the dual Hodge module $\mathbb{D}_{Y}(M)$ to be the mixed Hodge module satisfying
(3.2) $\mathbb{D}_{Y}(\mathrm{gr}_{k}^{W}M)=\mathrm{gr}_{-k}^{W}\mathbb{D}_{Y}(M)$
for every $k$. The underlying perverse sheaf $\mathrm{rat}~{}\mathbb{D}(M)$ is the Verdier dual of $\mathrm{rat}~{}M$, and the underlying $\mathscr{D}_{Y}$-module is the holonomic dual
$\mathbb{D}_{Y}(\mathcal{M})=R^{d}\mathrm{Hom}_{\mathscr{D}_{Y}}(\omega_{Y}\otimes_{\mathscr{O}_{Y}}\mathcal{M},\mathscr{D}_{Y}),$
which is compatible with the filtration $F_{\bullet}$ of item (ii) of §3.1.1, by [Sai88, Lemma 5.1.13]. In particular, if $M$ is a pure Hodge module of weight $w$, then $\mathbb{D}_{Y}(M)$ is a pure Hodge module of weight $-w$. In this case, $\mathbb{D}_{Y}(M)\cong M(w)$, where $M(w)$ is the Tate-twist of $M$ in degree $w$, obtained by twisting the perverse structure and Hodge filtration in the usual way.
The compatibility of the Hodge filtration with $\mathbb{D}_{Y}$ is due to Saito [Sai88, 2.4.3]. Specifically, there is an isomorphism
(3.3) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{M})),$
where again $\omega_{Y}^{\bullet}=\omega_{Y}[d]$. If $M\in\mathrm{HM}(Y,w)$, the isomorphism $\mathbb{D}_{Y}(M)\cong M(w)$ reduces (3.3) to
(3.4) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p-w}^{F}\mathrm{DR}(\mathcal{M}).$
3.1.3. The Restricted Hodge Module
Now let $Y=\mathbb{C}^{d}$ and $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$. For a generic hyperplane section $H$ of $Y$, $H$ intersects any Whitney strata adapted to the perverse sheaf $\mathrm{DR}(\mathcal{M})$ transversely. Therefore $H$ defines a non-characteristic hypersurface with respect to the left $\mathscr{D}_{Y}$-module $\mathcal{M}$, see [KS21, Definition 4.15] and [Sch16, §9]. Given such a hyperplane $H\subset Y$, we can construct a mixed Hodge module $M_{H}\in\mathrm{MHM}(H)$ by [Sai90, Lemma 2.25]. If $\iota_{H}:H\hookrightarrow Y$ is the inclusion, the underlying $\mathscr{D}_{H}$-module is
$\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}\mathcal{M}$
with filtration $F_{\bullet}\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}F_{\bullet}\mathcal{M}$, see [Sch16, Lemma 9.5]. The de Rham complex of $\mathcal{M}_{H}$ is
$\mathrm{DR}(\mathcal{M}_{H})=\iota_{H}^{-1}\mathrm{DR}(\mathcal{M})[-1].$
This give the data of a mixed Hodge module $M_{H}$ on $H$.
The restricted Hodge module will be important for many inductive arguments, as we have the following generalization of the conormal bundle sequence [Sch19, (13.3)]:
(3.5) $0\to N_{H|Y}^{*}\otimes_{\mathscr{O}_{H}}\mathrm{gr}_{p+1}^{F}\mathrm{DR}(\mathcal{M}_{H})\to\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M})\to\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}_{H})[1]\to 0,$
where $N_{H|Y}^{*}$ is the conormal bundle of the inclusion $\iota_{H}$.
3.2. The Intersection Hodge Module
Returning now to singularities, let $X$ be a reduced and (for convenience) irreducible complex analytic variety of dimension $n$, and let $\iota:X\hookrightarrow Y:=\mathbb{C}^{n+c}$ be a (local) closed embedding into the smooth open ball $Y$ of codimension $c$. We consider an object $\mathrm{IC}_{X}\in\mathrm{HM}(Y,n)$, called the intersection Hodge module, whose support is exactly $X$ [^2] More generally, we can consider the category of pure/mixed Hodge modules on $X$ by passing to local embeddings into smooth varieties: see [Sai90, §2]. By [Sai90, Thm. 3.21], the category $\mathrm{PVHS}_{\mathrm{gen}}(X,w)$, which is the direct limit of polarized variations of Hodge structures with quasi-unipotent local monodromies over Zariski open subsets $U\subset X\subset Y$, is equivalent to the subcategory of $\mathrm{HM}(Y,w)$ of pure Hodge modules with strict support $X$.
Definition 3.2.
Let $X$ be a complex analytic variety of dimension $n$ and $\iota:X\hookrightarrow Y$ a smooth embedding into a ball $Y=\mathbb{C}^{n+c}$ of codimension $c$. The intersection Hodge module $\mathrm{IC}_{X}=(\mathcal{IC}_{X},F,IC_{X})\in\mathrm{HM}(Y,n)$ is the unique Hodge module with support $X$ determined by the variation of Hodge structures $\mathbb{Q}_{U}[n]$, where $U=X_{\mathrm{reg}}$ is the regular locus of $X$.
We remark that the underlying perverse sheaf $IC_{X}$ of $\mathbb{Q}$-constructible sheaves computes the intersection cohomology groups $IH^{\bullet}(X,\mathbb{Q}):=\mathbb{H}^{\bullet-n}(X,IC_{X})$ of $X$. With this in mind, we get a Hodge-theoretic interpretation of the decomposition theorem.
Let $\pi:\widetilde{X}\to X$ be a resolution of singularities, and consider the induced morphism $f=\iota\circ\pi:\widetilde{X}\to Y$. By Saito’s direct image theorem [Sai88, §5.3], there are pure Hodge modules $M_{l}=(\mathcal{M}_{l},F_{\bullet},\mathrm{rat}~{}M_{l})$ supported on the singularities of $X$ and a decomposition
(3.6) $\mathbf{R}f_{*}\Omega_{\widetilde{X}}^{p}[n-p]\cong_{\mathrm{qis}}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})\oplus\bigoplus_{l\in\mathbb{Z}}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{M}_{l})[-l],$
see [KS21, (8.0.3)]. As the Hodge modules $M_{l}$ are supported on the singularities, they are torsion; this gives
(3.7) $f_{*}\Omega_{\widetilde{X}}^{p}\cong\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})$
by [KS21, Proposition 8.1].
3.3. The Logarithmic Hodge Module
Let $X$ be a reduced and irreducible complex analytic variety and $\iota:X\hookrightarrow Y$ a (local) closed embedding into a complex ball $Y=\mathbb{C}^{n+c}$ of codimension $c$. If $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with log-exceptional divisor $E$, recall from §2.1 that we assume an isomorphism $U:=X_{\mathrm{reg}}\cong\widetilde{X}\setminus E$. The sheaf $\mathbb{Q}_{\widetilde{X}\setminus E}$ defines a variation of pure Hodge structures; as a perverse sheaf, we can consider the pushforward $\mathbf{R}j_{*}\mathbb{Q}_{\widetilde{X}\setminus E}$, where $j:\tilde{X}\setminus E\hookrightarrow\widetilde{X}$ is the inclusion. By [Sai90, Thm. 3.27], this extends to a graded-polarizble mixed Hodge module on the smooth manifold $\widetilde{X}$. Specifically, the underlying $\mathscr{D}_{\widetilde{X}}$-module is the sheaf $\mathscr{O}_{\widetilde{X}}(*E)$ of meromorphic functions which are regular outside $E$. The filtered pieces $F_{p}\mathscr{D}_{\widetilde{X}}$ act naturally on $\mathscr{O}_{\widetilde{X}}(*E)$, inducing a good filtration on $\mathscr{O}_{\widetilde{X}}(*E)$. The de Rham complex is simply
[…]
By [Sai90, Proposition 3.11] (or classically), the inclusion $\Omega_{\widetilde{X}}^{\bullet}(\log E)[n]\hookrightarrow\mathrm{DR}(\mathscr{O}_{\widetilde{X}}(*E))$ is a filtered quasi-isomorphism:
$\Omega_{\widetilde{X}}^{p}(\log E)[n-p]\cong_{\mathrm{qis}}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathscr{O}_{\widetilde{X}}(*E)).$
Let $f=\iota\circ\pi:\widetilde{X}\to Y$. Saito’s direct image theorem says there is a family of mixed Hodge modules $\left\{N_{l}\right\}_{l\in\mathbb{Z}}$ whose rational perverse structure $\mathrm{rat}~{}N_{l}$ is simply the $l^{th}$ perverse cohomology sheaf of $\mathbf{R}f_{*}(j_{*}\mathbb{Q}_{\widetilde{X}\setminus E}[n])\cong\mathbf{R}j_{*}\mathbb{Q}_{U}[n]$. As such, the support of $N_{0}$ is $X$, the support of $N_{l}$ for $l\neq 0$ is contained in $X_{\mathrm{sing}}$, and $N_{0}$ has no non-trivial sub-objects supported on $X_{\mathrm{sing}}$ [KS21, Lemma 9.3].
Definition 3.3.
Let $X$ be a reduced and irreducible complex analytic variety of dimension $n$, let $\iota:X\to Y$ be a (local) closed embedding into $Y=\mathbb{C}^{n+c}$, let $\pi:\widetilde{X}\to X$ be a log-resolution of singularities with log-exceptional divisor $E$, let $j:\widetilde{X}\setminus E\hookrightarrow\widetilde{X}$ be the inclusion, and let $f=\iota\circ\pi:\widetilde{X}\to Y$.
The logarithmic mixed Hodge module $N_{0}\in\mathrm{MHM}(Y)$ is the unique mixed Hodge module supported on $X$ obtained from the push-forward $f_{*}(j_{*}\mathbb{Q}_{\widetilde{X}\setminus E})$ of perverse sheaves. We refer to the underlying $\mathscr{D}_{Y}$-module as $\mathcal{N}_{0}$.
By [KS21, Proposition 9.5], we get the following important relationship between logarithmic forms and the logarithmic Hodge module $N_{0}=(\mathcal{N}_{0},F_{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$:
(3.8) $f_{*}\Omega_{\widetilde{X}}^{p}(\log E)\cong\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0}).$
3.4. An Extension Criterion for Differentials on a Resolution of Singularities
Let $X$ be a normal complex analytic variety. By (3.7) and (3.8) the sheaves $\pi_{*}\Omega_{\widetilde{X}}^{p}$ and $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ associated to a (log-)resolution of singularities can be recovered from the intersection Hodge module $\mathrm{IC}_{X}$ and the logarithmic Hodge module $N_{0}$, respectively. By Proposition 1.5, the following gives a criterion to when the inclusions (1.1) and (1.2) are isomorphisms:
Proposition 3.4.
Let $X$ be a normal complex variety of dimension $n$, and let $X\subset Y$ be a (local) closed embedding into a smooth complex manifold $Y$.
  1. (i)
    Holomorphic extension holds in degree $p\geq 0$ if
    (3.9) $\dim\mathrm{Supp}~{}\mathscr{H}^{j+n-p}\mathrm{gr}_{-(n-p)}^{F}\mathrm{DR}(\mathcal{IC}_{X})\leq-(j+2)$
    for every $j\geq 0$, where $\mathcal{IC}_{X}$ is the $\mathscr{D}_{Y}$-module underlying the intersection Hodge module $\mathrm{IC}_{X}$.
  2. (ii)
    Logarithmic extension holds in degree $p\geq 0$ if
    (3.10) $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0}))\leq n-j-p-2$
    for every $j\geq 0$, where $\mathcal{N}_{0}$ is the $\mathscr{D}_{Y}$-module underlying the logarithmic Hodge module $N_{0}$.
Proof.
Rewrite (3.7) and (3.8) as
$f_{*}\Omega_{\widetilde{X}}^{p}\cong\mathscr{H}^{0}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})[p-n],\quad f_{*}\Omega_{\widetilde{X}}^{p}(\log E)\cong\mathscr{H}^{0}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0})[p-n].$
Proposition 1.5 says then that these global sections extend uniquely in codimension 2 if
$\dim\mathrm{Supp}~{}\mathscr{H}^{j+n-p}\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})[p-n],\omega_{Y}^{\bullet})\leq(-j+2)$
and
$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0})[p-n],\omega_{Y}^{\bullet})\leq-(j+2)$
hold, respectively. The first inequality is equivalent to (3.9) by duality for pure Hodge modules (3.4), since $\mathrm{IC}_{X}$ has weight $n$, and the second inequality is equivalent to (3.10) by duality for mixed Hodge modules (3.3). ∎
4. Logarithmic Extension for du Bois Singularities
4.1. Work of Kovács-Schwede-Smith
Let $X$ be a normal variety of dimension $n$ with singular locus $\Sigma$. It is well-known that holomorphic extension fails in degrees $p\geq\mathrm{codim}_{X}(\Sigma)$ for varieties with du Bois singularities: the affine cone of a smooth and projective Calabi-Yau variety will be strictly log-canonical and will fail holomorphic extension in degree $n$. On the other hand, logarithmic extension is known to hold in all degrees if we further assume $X$ is Cohen-Macaulay by [KSS10, Theorem 1.1] and [KS21, Theorem 1.5]. The key input is to use Proposition 2.3 to identify the sheaf $\pi_{*}\omega_{\widetilde{X}}(E)$ of logarithmic $n$-forms coming from a log-resolution $\pi:\widetilde{X}\to X$ with $\mathscr{H}^{0}\mathbb{D}(\underline{\Omega}_{X}^{0})$. We extend this proof with minor adjustments to the non-CM case.
4.2. Proof of Theorem 1.1
Lemma 4.1.
Let $X$ be a normal complex algebraic variety of dimension $n$. For any log-resolution of singularities $\pi:\widetilde{X}\to X$ with log-exceptional divisor $E$, then the natural map
$\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{0})\to\pi_{*}\omega_{\widetilde{X}}(E)$
is an isomorphism, where $\mathbb{D}_{X}$ is the Grothendieck duality functor on $X$.
Proof.
Let $X\subset Y$ be a (locally) closed embedding into a smooth complex manifold $Y$, and consider the singular locus $\Sigma\subset X\subset Y$ under this embedding. There is a distinguished triangle
$\mathbf{R}\mathscr{H}om_{\mathscr{O}_{X}}(\underline{\Omega}_{\Sigma}^{0},\omega_{X}^{\bullet})\to\mathbf{R}\mathscr{H}om_{\mathscr{O}_{X}}(\underline{\Omega}_{X}^{0},\omega_{X}^{\bullet})\to\mathbf{R}\pi_{*}\omega_{\widetilde{X}}(E)[n]\xrightarrow{+1}$
obtained by dualizing the triangle of Theorem 2.1(ix) for $p=0$. By [KSS10, Corollary 3.7], $R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(\underline{\Omega}_{\Sigma}^{0},\omega_{X}^{\bullet})=0$ for $j<-\dim\Sigma$. Specifically, there is a spectral sequence
$E_{2}^{p,q}=R^{p}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathscr{H}^{-q}\underline{\Omega}_{\Sigma}^{0},\omega_{Y}^{\bullet})\Rightarrow R^{p+q}\mathscr{H}om_{\mathscr{O}_{Y}}(\underline{\Omega}_{\Sigma}^{0},\omega_{Y}^{\bullet}).$
Note that $\mathrm{Supp}~{}E_{2}^{p,q}\neq 0$ if $j=-\dim\Sigma,...,0$, whence $\dim\mathrm{Supp}~{}E_{2}^{p,q}=0$ if $p<\dim Z-q$ by Theorem 2.1(iv). Thus $E_{\infty}^{p,q}=0$ in this range, which gives the desired vanishing on $X$. ∎
Proof of Theorem 1.1. By assumption, $\underline{\Omega}_{X}^{0}$ has no higher cohomology, and $\dim\mathrm{Supp}~{}\mathscr{O}_{X}<n-2$ since $X$ is normal. Therefore logarithmic extension holds in degree $n$ by Lemma 4.1 and Proposition 1.5. By [KS21, Theorem 1.5], logarithmic extension holds in all degrees $0\leq p\leq n$. ∎
5. Holomorphic Extension for du Bois Singularities
5.1. Consequences of Logarithmic Extension
Let $X$ be a normal variety of dimension $n$ and $\pi:\widetilde{X}\to X$ a log-resolution of singularities. For dimension reasons and (3.8), there is a quasi-isomorphism
$\pi_{*}\omega_{\widetilde{X}}(E)\cong\mathrm{gr}_{-n}^{F}\mathrm{DR}(\mathcal{N}_{0}).$
If $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive, we get a family of inequalities
$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))\leq-(j+p+2)$
for $p+j\geq-n+1$, see [KS21, Proposition 9.10]. This leads to an important vanishing as a consequence of Theorem 1.1.
Corollary 5.1.
Let $X$ be a normal complex analytic variety of dimension $n$ with at worst du Bois singularities. Then
$\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))=0,$
where $X\subset Y$ is a (locally) closed embedding into a smooth manifold $Y$ and $\mathcal{N}_{0}$ is the $\mathscr{D}_{Y}$-module underlying the logarithmic mixed Hodge module.
This is the main component to beginning the induction for the holomorphic extension property for du Bois singularities, using the relationship between the weight filtration of the logarithmic Hodge module and the intersection Hodge module.
Proposition 5.2.
Let $X$ be a normal variety of dimension $n$ with at worst du Bois singularities, and let $\mathrm{IC}_{X}=(\mathcal{I}C_{X},F_{\bullet},IC_{X})$ be the intersection Hodge module. Then
$\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathcal{IC}_{X})=0.$
Proof.
Let $X\subset Y$ be a (locally) closed embedding into a smooth complex manifold $Y$. Let $N_{0}=(\mathcal{N}_{0},F_{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$ be the logarithmic mixed Hodge module. By construction, $W_{n}\mathbb{D}_{Y}(\mathcal{N}_{0})=\mathbb{D}_{Y}(W_{n}\mathcal{N}_{0})$. Since $\mathrm{gr}_{-1}^{F}\mathrm{DR}(-)$ is an exact functor, there is an exact sequence
$\begin{split}\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))&\to\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}(\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))\\ &\to\mathscr{H}^{1}\mathrm{gr}_{-1}^{F}\mathrm{DR}(W_{n-1}\mathbb{D}(\mathcal{N}_{0})(-n)).\end{split}$
The last term must be zero for degree reasons, and so by Corollary 5.1 we get the additonal vanishing $\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))=0$. By (3.3), we have an isomorphism $\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n)\cong\mathrm{gr}_{n}^{W}\mathcal{N}_{0}$. The claim follows from the isomorphism
$\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathrm{gr}_{n}^{W}\mathcal{N}_{0})\cong_{\mathrm{qis}}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathcal{IC_{X}}),$
see [KS21, Proposition 9.8]. ∎
5.2. Proof of Theorem 1.2
Let $X$ be a normal variety of dimension $n$ with du Bois singularities. Recall from Proposition 3.4 that holomorphic extension holds in degree $p$ if
(5.1) $\mathrm{dim}~{}\mathrm{Supp}~{}\mathscr{H}^{j+n-p}\mathrm{gr}_{-(n-p)}^{F}\mathrm{DR}(\mathcal{IC}_{X})\leq-(j+2)$
Note that this condition vacuously holds for isolated singularities except possibly when $(p,j)=(n-1,0)$, which is covered by Proposition 5.2. Therefore, Theorem 1.2 holds for isolated singularities (compare to §1.3).
Proof of Theorem 1.2. We have demonstrated (5.1) for normal surfaces with du Bois singularities. We proceed then by induction on the pair $(n,\dim\Sigma)$, ordered lexicographically. If $X\subset Y$ a closed embedding into a smooth ball $Y=\mathbb{C}^{n+c}$, let $H\subset Y$ be a general hyperplane section and consider the exact sequence of complexes
$0\to N_{H|Y}^{*}\otimes_{\mathscr{O}_{H}}\mathrm{gr}_{p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X}|_{H})\to\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{IC}_{X})\to\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{IC}_{X}|_{H})[1]\to 0$
of (3.5). Since $H$ is generic, the restricted Hodge module $\mathrm{IC}_{X}|_{H}=\mathrm{IC}_{X\cap H}$ since we can assume that $(X\cap H)_{\mathrm{reg}}=X_{\mathrm{reg}}\cap H$ (see Definition 3.2). Since $H$ is du Bois by 2.1(vi), then (5.1) and induction imply
$\dim\mathrm{Supp}~{}\mathscr{O}_{H}\otimes\mathscr{H}^{j+(n-1)-p}\mathrm{gr}_{-(n-p)}^{F}\mathrm{DR}(\mathcal{IC}_{X})\leq-(j+2)$
for $p=\mathrm{codim}_{X}(\Sigma)$. Therefore
(5.2) $\dim\mathrm{Supp}~{}\mathscr{H}^{j+n-p}\mathrm{gr}_{-(n-p)}^{F}\mathrm{DR}(\mathcal{IC}_{X})\leq-(j+2),$
except possibly in the case $(p,j)=(1,0)$. But this is exactly Proposition 5.2. ∎
5.3. Log Forms to Holomorphic Forms
Suppose now that $X$ is an arbitrary normal complex variety with possibly non-du Bois singularities, and let $\Sigma$ be the singular locus. It is known that holomorphic extension can fail for $p\geq\mathrm{codim}_{X}(\Sigma)-1$. We consider then the inclusion (1.4) and give a weakening of Theorem 1.2 to arbitrary singularities. We remark that this theorem is optimal even for du Bois singularities.
Proof of Theorem 1.7. Let $\mathrm{IC}_{X}$ and $N_{0}$ be the intersection and logarithmic Hodge modules of the normal variety $X$. We note there is a canonical morphism of (mixed) Hodge modules $M_{X}\to N_{0}$ obtained by the isomorphism $W_{n}N_{0}\cong\mathrm{IC}_{X}$ [KS21, Proof of (9.8.1)]. The problem is local, so let $X\subset Y$ be a closed embedding of $X$ into a ball $Y\cong\mathbb{C}^{n+c}$, and let $H\subset Y$ be a generic hyperplane section. For each $p$, there is a commutative diagram of exact sequences
[…]
coming from the exact triangle associated to the restricted Hodge module sequence (3.5). Again, $\mathcal{IC}_{X\cap H}$ (resp. $\mathcal{N}_{0}|_{H}$) is the intersection $\mathscr{D}_{Y}$-module (resp. logarithmic $\mathscr{D}_{Y}$-module) of $X\cap H$. By (3.7) and (3.8), the first three rows of the commutative diagram agree with (1.5).
We note that if $X$ has isolated singularities, the claim is known to hold [vSS85, Thm. (1.3)]. Continuing by induction, consider the above diagram for $p=\mathrm{codim}_{X}(\Sigma)-1$. The first horizontal morphism is an isomorphism by Flenner’s criterion, and the third horizontal morphism is an isomorphism by induction. For this $p$, we also have
$\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})=\mathscr{H}^{-\dim\Sigma}\mathrm{gr}_{-((n-1)-\dim\Sigma)}\mathrm{DR}(\mathcal{IC}_{X\cap H}).$
For dimension reasons, this term vanishes. Therefore the morphism
$\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$
is an isomorphism by the 4-lemma. The theorem follows since $H$ was chosen to be general. ∎
This gives a different proof of Theorem 1.2 for du Bois singularities by Theorem 1.1. We emphasize that this proof is weaker, however, since it does not imply support condition (5.1).
6. Holomorphic Extension for (Weakly) $k$-du Bois Singularities
6.1. Rational v.s. Weakly $k$-du Bois Singularities
A fundamental aspect of the Hodge theory of singularities is the relationship between the du Bois property and the singularities of the MMP: du Bois singularities are very close to log-canonical singularities, as both agree in the normal, quasi-Gorenstein case. The gap between rational and du Bois singularities is much larger. Recall by Kempf’s criterion that a variety has rational singularities if and only if $X$ is Cohen-Macaulay $\pi_{*}\omega_{\tilde{X}}$ is reflexive. We have already seen a du Bois singularity failing holomorphic extension, but they can also fail the Cohen-Macaulay property: the affine cone of a compact hyperkähler manifold in dimension $\geq 4$ is du Bois but not Cohen-Macaulay.
Therefore, it seems interesting to ask how closely (weakly) $k$-du Bois singularities are from having rational singularities. Even for $k>0$, there is a disconnect. The following affine cone examples were described in [Tig23], see also [SVV23, §7.5].
Example 6.1.
Let $Y$ be a projective K3 surface. For any ample bundle $L$, the affine cone $X$ is Cohen-Macaulay and du Bois by Kodaira vanishing. Suppose further that $Y$ has Picard rank 1 and the degree $L^{2}\geq 24$. Then the affine cone is also weakly 1-du Bois, and $\mathscr{H}^{0}\underline{\Omega}_{X}^{1}$ is reflexive. Indeed, it is sufficient to show that $H^{1}(Y,\Omega_{Y}^{1}\otimes L^{m})=0$ for every $m>0$. This follows from [Tot20, Theorem 3.2 and Theorem 3.5]. The first theorem states that, under these assumptions, the vanishing holds for $H^{1}(Y,\Omega_{Y}^{1}\otimes L)=0$, considering the pair $(Y,L)$ as a polarized K3 surface. The second theorem verifies the vanishing $H^{1}(Y,\Omega_{Y}^{1}\otimes L^{m})=0$ for $m>1$ assuming the first vanishing. Since $H^{0}(Y,\Omega_{Y}^{1})=0$, this implies that $X$ is weakly 1-du Bois. But $H^{0}(Y,\omega_{Y})\cong\mathbb{C}$, and so $X$ does not have rational singularities. By Theorem 1.2, $\pi_{*}\Omega_{\widetilde{X}}^{2}$ is reflexive for any resolution of singularities.
Example 6.2.
Here is an example of an affine cone singularity which is rational but is not $k$-du Bois for some $11\leq k\leq n-$ 1. The example is essentially given in [BTLM97, §4.1] and is related to the failure of Bott vanishing for non-projective spaces. Let $Y\subset\mathbb{P}^{4}$ be a smooth quadric hypersurface. Let $L=\mathscr{O}_{Y}(1)$. Then the affine cone $X$ has rational singularities since $Y$ is a Fano variety. The cone is also weakly 1-du Bois. To see this [^3] Since $Y$ is a hypersurface, then so will the affine cone $X$. Therefore this follows from [MOPW21], we consider the cohomology sequence of
$0\to\mathscr{O}_{Y}(2+m)\to\Omega_{\mathbb{P}^{4}}^{1}\otimes\mathscr{O}(m)\otimes\mathscr{O}_{Y}\to\Omega_{Y}^{1}\otimes\mathscr{O}_{Y}(m)\to 0$
The vanishing $H^{i}(Y,\Omega_{Y}^{1}\otimes\mathscr{O}_{Y}(m))=0$ follows then by Bott vanishing on $\mathbb{P}^{4}$ and the rationality of $(X,v)$. On the other hand, $H^{1}(Y,\Omega_{Y}^{2}\otimes\mathscr{O}_{Y}(1))\cong\mathbb{C}$. This implies that $X$ is not weakly $2$-du Bois.
In summary: weakly $k$-du Bois singularities need not have rational singularities, nor are rational singularities weakly $k$-du Bois for all $k$. What is interesting about Example 6.1 is that the 3-fold singularity is weakly $3$-du Bois, and yet holomorphic extension does not hold in degree 3. This is because $\mathscr{H}^{0}\underline{\Omega}_{X}^{2}$ is also not reflexive in this case, a necessary condition for holomorphic extension to hold for $k$-du Bois singularities.
6.2. Isolated Singularities and Depth
To further highlight what Example 6.1 tells us, we consider the relationship between $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$, $\pi_{*}\Omega_{\widetilde{X}}^{p}$, and $\mathrm{depth}(\mathscr{O}_{X})$. Let $X$ be a normal variety with isolated singularities. If $X$ is weakly $k$-du Bois, then Theorem 2.1(vii) implies a short exact sequence
$0\to\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\mathscr{\pi}_{*}\Omega_{\tilde{X}}^{p}\to\pi_{*}\underline{\Omega}_{E}^{p}\to 0$
for $0\leq p\leq k$. An immediate consequence of Theorem 2.1(iii) is that $\pi_{*}\underline{\Omega}_{E}^{p}=0$ if $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive. We note that $\underline{\Omega}_{E}^{p}=\Omega_{E}^{p}/\mathrm{tor}$ is just a sheaf; if $E$ is an snc divisor, then $H^{0}(E,\underline{\Omega}_{E}^{p})$ vanishes by Hodge theory if $H^{p}(E,\mathscr{O}_{E})=0$. This for example holds if $\mathrm{depth}(\mathscr{O}_{X})\geq p+2$ (compare with [MP22, Theorem G]).
Note that even if we assume $X$ is Cohen-Macaulay that we do not get the vanishing of $\pi_{*}\underline{\Omega}_{E}^{\dim X-1}$: this is precisely what happens in Example 6.1. The reflexivity of $\mathscr{H}^{0}\underline{\Omega}_{X}^{\dim X-1}$ appears independent of any condition on the depth, similar to the sheaf $\pi_{*}\omega_{\widetilde{X}}$.
6.3. A criterion for Holomorphic Extension for Weakly $k$-du Bois Singularities
Theorem 6.3.
Suppose $X$ is a normal complex algebraic variety of dimension $n$. If $X$ is weakly $k$-du Bois and $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then holomorphic extension holds in degree $k+1$.
Proof.
By Proposition 1.3 and the quasi-isomorphism $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{k}$, the assumption on $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ implies $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathbb{D}_{X}(\underline{\Omega}_{X}^{k})\leq-(j+2)$ for $j\geq-n+1$. By looking exact triangle of Theorem 2.1(ix), dualizing, and taking cohomology, we get the vanishing $R^{n-1}\pi_{*}\Omega_{\widetilde{X}}^{n-k}(\log E)=0$. By the residue exact sequence, $R^{n-1}\pi_{*}\Omega_{E_{i}}^{n-k-1}=0$ for each smooth component $E_{i}$ of $E$. Since the problem is local, we can shrink $X$ and assume that $H^{n-1}(E_{i},\Omega_{E_{i}}^{n-k-1})=0$. The $E_{i}$ are smooth and projective (see §2.1), so $H^{0}(E_{i},\Omega_{E_{i}}^{k})=0$, too. This implies the inclusion $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{k+1})\hookrightarrow H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{k+1}(\log E))$ is an isomorphism. Since $X$ is du Bois, the claim follows from Theorem 1.2. ∎
Corollary 6.4.
Let $X$ be a normal complex variety with weakly $k$-du Bois singularities for $k\geq\mathrm{codim}_{X}(\Sigma)-1$, where $\Sigma$ is the singular locus of $X$. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for each $0\leq p\leq k$, then
(6.1) $\dim\mathrm{Supp}~{}R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}\leq n-j-2$
for each $j\leq k$.
Proof.
For isolated singularities, the assumption on $k$ implies $\pi_{*}\omega_{\widetilde{X}}$ is reflexive by Theorem 6.3. Therefore the claim (6.1) holds by Proposition 1.3 and Grauert-Rimenschneider vanishing. More generally, let $X\subset Y$ be a locally closed embedding into a smooth manifold $Y$ and let $H$ be a general hyperplane section of $Y$. By Theorem 2.1(vi), $X\cap H$ has weakly $k$-du Bois singularities, and it is clear that $\mathscr{H}^{0}\underline{\Omega}_{X\cap H}^{p}$ is reflexive whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive. Therefore induction and the isomorphism
$\mathscr{O}_{H}\otimes\mathbf{R}\pi_{*}\mathscr{O}_{\tilde{X}}\cong\mathbf{R}(\pi|_{H})_{*}\mathscr{O}_{\widetilde{X\cap H}}[1],$
where $\pi|_{H}$ is the induced resolution of singularities of $X\cap H$ from $\pi:\widetilde{X}\to X$, imply the claim except in the case $j=n-1$; but this is only relevant when $X$ is $(n-1)$-du Bois and so follows from Theorem 6.3. ∎
6.4. A Remark on the Functorial Pullback Morphism
Corollary 6.4 is optimal for weakly $k$-du Bois singularities. If $X$ is Cohen-Macaulay, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq\mathrm{codim}_{X}(\Sigma)-2$. This follows since:
  • $R^{p}\pi_{*}\mathscr{O}_{\widetilde{X}}=0$ for $p\leq\mathrm{codim}_{X}(\Sigma)-2$ [Kov99, Lemma 3.3], and
  • The holomorphic extension property holds for $p\leq\mathrm{codim}_{X}(\Sigma)-2$ by Flenner’s criterion.
The inclusion $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\hookrightarrow\pi_{*}\Omega_{\widetilde{X}}^{p}$ is therefore an isomorphism due to an idea of Kebekus-Schnell on the existence of functorial pullback morphisms.
Recall that if $f:Z\to X$ is a morphism of complex spaces, there is a functorial pullback morphism $f^{*}\Omega_{X}^{p}\to\Omega_{Z}^{p}$ between the sheaves of Kähler $p$-forms. In general, this pullback does not extend to the sheaves of reflexive differentials. Work of Kebekus [Keb13] describes a process of constructing a natural reflexive pullback morphism which agrees with the Kähler pullback morphism on the regular locus. The pullback morphism was originally constructed for morphisms of algebraic varieties with at worst klt singularities but was extended to arbitrary complex spaces with rational singularities [KS21, §14]. There are two major inputs for a pullback morphism $f^{*}\Omega_{X}^{[p]}\to\Omega_{Z}^{[p]}$ to exist in degree $p$, which hold for rational singularities:
  • The varieties $X$ and $Z$ satisfy the holomorphic extension property in degree $p$ [KS21, Corollary 1.8]
  • If $E_{t}=\pi^{-1}(t)$ is the fiber of a resolution of singularities $\pi:\tilde{X}\to X$, then $H^{0}(E_{t},\Omega_{E_{t}}^{p}/\mathrm{tor})=0$ [Nam01, Lemma 1.2].
Recall from §6.2 that the second condition holds for isolated weakly $k$-du Bois singularities whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq k$. The first condition also holds in degree $p=k+1$ by Theorem 6.3. Since $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ agrees with the sheafification of $\Omega_{X}^{p}$ in the $h$-topology [HJ14], $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive whenever these conditions for functorial pullback hold. The following corollary of Theorem 6.3 is therefore immediate.
Corollary 6.5.
Let $X$ be a normal complex variety with singular locus $\Sigma$, and let $\pi:\widetilde{X}\to X$ be a log-resolution of singularities. Suppose $X$ is weakly $k$-du Bois for $k\geq\mathrm{codim}_{X}(\Sigma)-1$ and that $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for some $p\leq k$. Then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p+1}$ is reflexive if and only if $\pi_{*}\Omega_{E_{i}}^{p+1}=0$ for each smooth component $E_{i}$ of the log-exceptional divisor $E$ of $\pi$.
6.5. A Remark on the Reflexivity of $\Omega_{X}^{p}$
Let $X$ be a normal complex algebraic variety. We wish to consider the $k$-du Bois and $k$-rational definitions defined in [MOPW21], [JKSY21], [MP22], [FL22a], and [FL22b]. These papers consider lci singularities, in which case the $k$-du Bois properties implies $\Omega_{X}^{p}$ is reflexive for every $p$; this is particularly special for lci singularities and restricts the codimension of the singular locus. For instance, if $Y$ is a variety with lci singularities, $\Omega_{Y}^{1}$ is reflexive if and only if $Y$ is smooth in codimension 3 [Kun86].
Understanding what happens when $\Omega_{X}^{p}$ is reflexive seems to be a difficult problem. Instead, we consider what happens when $X$ is $k$-rational for large $k$.
Theorem 6.6.
If $X$ is a complex algebraic variety with gorenstein $k$-rational singularities for $k>\frac{1}{2}\dim X$, then $X$ is smooth.
Proof.
Let $x\in X$. For $p\leq k$, consider the stalk $\Omega_{X,x}^{p}$. Then $\mathrm{Ext}^{j}_{\mathscr{O}_{X,x}}(\Omega_{X,x}^{p},\omega_{X,x})\cong\mathscr{H}^{j}\mathbb{D}_{X}(\underline{\Omega}_{X}^{p})_{x}=0$ for $j>0$. This implies $\Omega_{X,x}^{p}$ is projective as an $\mathscr{O}_{X,x}$-module, and the same holds for $\mathscr{F}_{p}:=\Omega_{X,x}^{p}\otimes(\omega_{X}^{*})_{x}$ since $X$ is gorenstein. There is then a trace operator
$\mathscr{F}_{p}\otimes(\mathscr{F}_{p})^{*}\to\mathscr{O}_{X,x},$
and so any minimal generating sets of $\mathscr{F}_{p}$ and $(\mathscr{F}_{p})^{*}$ have the same cardinality, i.e., $\mathrm{rk}~{}\mathscr{F}_{p}=\mathrm{rk}~{}(\mathscr{F}_{p})^{*}$.
Let $n=\dim X$. By our assumption on $k$, there is a $p<\frac{1}{2}n$ such that $\Omega_{X}^{p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ and $\Omega_{X}^{n-p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{p})$. Then
$\Omega_{X}^{n-p}\cong\mathscr{H}om(\Omega_{X}^{p},\omega_{X})\cong\mathscr{H}om(\Omega_{X}^{p}\otimes\omega_{X}^{*},\mathscr{O}_{X}).$
Using again the fact that $X$ is Gorenstein, the preceding argument says that $\mathrm{rk}~{}\Omega_{X,x}^{p}=\mathrm{rk}~{}\Omega_{X,x}^{n-p}$. But $\mathrm{rk}~{}\Omega_{X,x}^{r}=\binom{e_{x}}{r}$ for any $r$, where $e_{x}$ is the embedding dimension of the singularity $(X,x)$ [Gra15, §4]. Since $e_{x}\geq n$, the equality $\binom{e_{x}}{p}=\binom{e_{x}}{n-p}$ implies that $e_{x}=n$. Therefore $X$ is smooth. ∎
Corollary 6.7.
If $X$ is a normal gorenstein variety with singular locus $\Sigma$. If $X$ is $k$-rational with $k>\frac{1}{2}\mathrm{codim}_{X}(\Sigma)$, then $X$ is smooth.
Proof.
The case of isolated singularities is covered by Theorem 6.6. The more general fact follows since a general hyperplane section of a $k$-rational singularity is $k$-rational [SVV23, Theorem A]. ∎
We remark that Corollary 6.7 was observed for the $k$-du Bois property in the case of hypersurface singularities [MOPW21, Lemma 2.2], which is equivalent to being $(k-1)$-rational [FL22b, Corollary 1.9]. We can recover this result for isolated singularities by the Proof of Theorem 6.6, since we only require:
  • $X$ is $k$-rational for some $k\leq\frac{1}{2}\dim X$,
  • $\Omega_{X}^{n-k}$ is reflexive.
While $\Omega_{X}^{n-k}$ may not be reflexive if $X$ is only $k$-du Bois, this always holds for hypersurface singularities.
7. Hodge Theory of Proper Varieties with Weakly $k$-Rational Singularities
7.1. Mixed Hodge Theory for MMP Singularities
Let $X$ be a complex algebraic variety. In his thesis, Deligne defined the notion of a mixed Hodge structure and constructed a canonical mixed Hodge structure on the cohomology of any complex algebraic variety [Del71], [Del74]. Specfically, there is an increasing weight filtration $W^{\bullet}$ on $H^{k}(X,\mathbb{Q})$ and a decreasing filtration $F^{\bullet}$ on $H^{k}(X,\mathbb{C})$ which descends to a Hodge filtration on $\mathrm{gr}_{p}^{W}H^{k}(X,\mathbb{Q})\otimes\mathbb{C}$ for each $k$. The du Bois complex is a byproduct of this construction: namely, du Bois details in [DB81] that Deligne’s construction produces a complex $\underline{\Omega}_{X}^{\bullet}$ for any algebraic variety $X$ which generates the Hodge filtration under the spectral sequence
$E_{1}^{p,q}=\mathbb{H}^{q}(X,\underline{\Omega}_{X}^{p})\Rightarrow H^{p+q}(X,\mathbb{C})$
when $X$ is proper.
It is often useful to understand when the mixed Hodge structure on $H^{k}(X,\mathbb{Q})$ is pure when studying the global moduli of singularities. For low degree, this is usually understood by looking at the pullback morphism $\pi^{*}:\pi^{*}H^{k}(X,\mathbb{Q})\to H^{k}(\widetilde{X},\mathbb{Q})$ associated to a resolution of singularities. By Leray, this map is injective if $k=1$ when $X$ is normal. For $k=2$, this map is again injective by Leray when $X$ has rational singularities.
The obvious question to ask is how does the (weakly) $k$-du Bois property affect the mixed Hodge theory on the cohomology $H^{*}(X,\mathbb{Q})$ of a projective variety. Surprisingly, the vanishing cohomology of the $\underline{\Omega}_{X}^{p}$ has little control over the weight filtration:
Example 7.1.
  1. (1)
    Let $X$ be a projective curve. For dimension reasons, $X$ is weakly $1$-du Bois, but $H^{1}(X,\mathbb{Q})$ will not carry a pure Hodge structure. For an explicit example, consider the nodal elliptic curve.
  2. (2)
    Let $X$ be a projective hypersurface of dimension 3 with ordinary double points. Note that $X$ has rational singularities and is $1$-du Bois by [FL22b, Corollary 1.9]. Therefore, $H^{2}(X,\mathbb{Q})$ carries a pure Hodge structure by the preceding discussion, but $H^{3}(X,\mathbb{Q})$ need not carry a pure Hodge structure
  3. (3)
    Let $Y$ be a projective hyperkähler 4-fold manifold. Suppose there exists a birational contraction $\phi:Y\to X$ of a Lagrangian submanifold $L\cong\mathbb{P}^{2}\subset Y$ to a point. Then $X$ is not weakly $k$-du Bois for $k>1$, but $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure for each $k$.
Moreover, $H^{k}(X,\mathbb{Q})$ will carry a pure Hodge structure for large $k$ for topological reasons: if $X$ is projective and has isolated singularities, then $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure of weight $k$ for $k>n$. What is interesting in this case is that dual $H^{k}(X,\mathbb{Q}(-n))^{*}$ carries a pure Hodge structure of weight $2n-k$. By Poincaré duality, this group is $H^{2n-k}(X_{\mathrm{reg}},\mathbb{Q})$, and $H^{2n-k}(X,\mathbb{Q})$ carries a pure Hodge structure if and only if $H^{2n-k}(X,\mathbb{Q})\to H^{2n-k}(X_{\mathrm{reg}},\mathbb{Q})$ is injective.
More generally, we can consider the weight filtrations for the mixed Hodge structures on $H^{k}(X,\mathbb{Q})$ and $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$, respectively. Since $X$ is proper, the weight filtration $W^{\bullet}$ truncates to $H^{k}(X,\mathbb{Q})$ for each $k$. The weight filtration on the dual $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$ therefore is supported in higher weights (compare this to the cohomology $H^{k}(X_{\mathrm{reg}},\mathbb{Q})$), and these groups are related to the Grothendieck duals $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ of the du Bois complex. It therefore seems better to consider how the $k$-rational property affects the weight filtration.
7.2. Purity for $k$-Rational Singularities
Theorem 7.2.
If $X$ is a projective variety of dimension $n$ with weakly $k$-rational singularities, then the canonical mixed Hodge structure on $H^{k}(X,\mathbb{Q})$ is pure of weight $k$.
Proof.
By [SVV23, Theorem B], we have
$\Omega_{X}^{[p]}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$
for every $0\leq p\leq k$.
On the one hand, the spectral sequence
$E_{1}^{p,q}=\mathbb{H}^{q}(X,\underline{\Omega}_{X}^{p})\Rightarrow H^{p+q}(X,\mathbb{C})$
generates the Hodge filtration. Note that the weight filtration $W_{\mathbb{C}}^{\bullet}$ is supported in lower weights, since $X$ is proper.
On the other hand, $\mathbb{H}^{q}(X,\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p}))\cong\mathrm{Hom}_{\mathscr{O}_{\mathrm{pt}}}(\mathbb{H}^{n-q}(X,\underline{\Omega}_{X}^{n-p}),\mathscr{O}_{\mathrm{pt}})$ by duality. Therefore $E_{1}^{p,q}$ generates the Hodge filtration on the dual mixed Hodge structure $H^{n-k}(X,\mathbb{C})^{*}$ for $k=p+q$. The weight filtration on this mixed Hodge structure is supported in higher weights. Therefore $H^{k}(X,\mathbb{Q})$ must be a pure Hodge structure. ∎
Corollary 7.3.
If $X$ is a projective variety of dimension $n$ with weakly $k$-rational singularities, then there is a non-canonical decomposition
$H^{k}(X,\mathbb{C})=\bigoplus_{p+q=k}H^{q}(X,\pi_{*}\Omega_{\tilde{X}}^{p})$
induced by the Hodge filtration.
We remark that Theorem 7.2 implies something stronger than the cohomology $H^{k}(X,\mathbb{Q})$ carrying a pure Hodge structure. For instance if $k<\mathrm{codim}_{X}(\Sigma)$, where $\Sigma$ is the singular locus, the above proof shows that $H^{k}(X,\mathbb{Q})\to H^{k}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism. If $X$ has rational singularities, this morphism is always injective for $k=2$. If $X$ is a 3-fold with isolated singularities, then $H^{2}(X,\mathbb{Q})\to H^{2}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism if and only if Poincaré duality holds: therefore the defect $\sigma(X)$ of $X$ must be 0 [Kaw88, p. 97], and $X$ is $\mathbb{Q}$-factorial. This extends to higher dimensions in special cases, see for example [Tig22, Proposition 2.17].

### v3

@@
-The Holomorphic Extension Property for $k$-du Bois Singularities
+The Holomorphic Extension Property for Higher Du Bois Singularities
 Abstract.
@@
 We show that the holomorphic extension property holds in degree $p<\mathrm{codim}_{X}(X_{\mathrm{sing}})$ when $X$ has Du Bois singularities, giving an improvement on Flenner’s criterion for arbitrary singularities.
-As an application, we study the $k$-du Bois definition from the perspective of holomorphic extension and compare how different restrictions on $\mathscr{H}^{0}(\underline{\Omega}_{X}^{p})$ affect the singularities of $X$, where $\underline{\Omega}_{X}^{p}$ is the $p^{th}$-graded piece of the du Bois complex.
+As an application, we study the $m$-Du Bois definition from the perspective of holomorphic extension and compare how different restrictions on $\mathscr{H}^{0}(\underline{\Omega}_{X}^{p})$ affect the singularities of $X$, where $\underline{\Omega}_{X}^{p}$ is the $p^{th}$-graded piece of the Du Bois complex.
 1. Introduction
@@
 Moreover, if holomorphic extension holds in degree $p$, then holomorphic extension holds in degree $k$ for every $0\leq k\leq p$ [KS21, Theorem 1.4].
-For $p\geq\mathrm{codim}_{X}(\Sigma)-1$, extension is less topological and (1.1) can be strict: by definition, strictly log-canonical singularities fail holomorphic extension in degree $\dim X$.
+For $p\geq\mathrm{codim}_{X}(\Sigma)-1$, extension is less topological and (1.1) can be strict: by definition, Gorenstein log-canonical singularities fail holomorphic extension in degree $\dim X$.
 Instead, it is better to consider a variant of the inclusion (1.1) which allows for logarithmic poles.
@@
 (1.3) $\mathscr{O}_{X}\to\underline{\Omega}_{X}^{0}$
-is a quasi-isomorphism, where $\underline{\Omega}_{X}^{0}:=\mathrm{gr}_{F}^{0}\underline{\Omega}_{X}^{\bullet}$ is the $0^{th}$-graded piece of the du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$, an object in the category of filtered complexes of $\mathscr{O}_{X}$-modules generalizing the holomorphic de Rham complex of smooth algebraic varieties.
-du Bois singularities are an important in the Hodge theory of singularities, as they arise in the construction of Deligne’s mixed Hodge structure [Del74] on the cohomology of algebraic varieties.
+is a quasi-isomorphism, where $\underline{\Omega}_{X}^{0}:=\mathrm{gr}^{F}_{0}\underline{\Omega}_{X}^{\bullet}$ is the $0^{th}$-graded piece of the Du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$, an object in the category of filtered complexes of $\mathscr{O}_{X}$-modules generalizing the holomorphic de Rham complex of smooth algebraic varieties.
+Du Bois singularities are an important class of singularities in the Hodge theory of singularities, as they arise in the construction of Deligne’s mixed Hodge structure [Del74] on the cohomology of algebraic varieties.
 They also play a fundamental role in the minimal model program, as log-canonical and rational singularities are Du Bois, see [KK10, Theorem 1.4] [Kov99, Corollary 2.6], [Sai00, 5.4.
@@
 If $X$ is Du Bois and Cohen-Macaulay, the answer to Question (ii) is already known: if $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, then the sheaf $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive by [KSS10, Theorem 1.1] [^1] In fact, a Cohen-Macaulay variety is Du Bois if and only if $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive, giving a natural generalization of Kempf’s criterion for rational singularities., and so $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is reflexive for all $0\leq p\leq\dim X$ [KS21, Theorem 1.5].
-We can extend the ideas of [KSS10] to give an analytic proof of logarithmic extension for normal du Bois singularities, giving a positive answer to Question (ii):
+This gives a short proof of a result of Kovács-Graf.
 Theorem 1.1.
-If $X$ is a normal complex algebraic variety with at worst du Bois singularities, then logarithmic extension holds in all degrees $0\leq p\leq\dim X$: for any logarithmic resolution of singularities $\pi:\widetilde{X}\to X$, the inclusion $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\hookrightarrow\Omega_{X}^{[p]}$ is an isomorphism.
+[GK14b, Theorem 4.1] If $X$ is a normal complex algebraic variety with at worst Du Bois singularities, then logarithmic extension holds in all degrees $0\leq p\leq\dim X$: for any logarithmic resolution of singularities $\pi:\widetilde{X}\to X$, the inclusion $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\hookrightarrow\Omega_{X}^{[p]}$ is an isomorphism.
 Theorem 1.1 is key to understanding for which $p$ holomorphic extension holds. Our main result is an extension of Flenner’s criterion to Du Bois singularities and an optimal answer to Question (i).
@@
 (1.4) $\pi_{*}\Omega_{\widetilde{X}}^{p}\hookrightarrow\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$
-is an isomorphism for $p=1$, where $\pi:\widetilde{X}\to X$ is a log-resolution of singularities, and showing $\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)$ is reflexive [GK14].
+is an isomorphism for $p=1$, where $\pi:\widetilde{X}\to X$ is a log-resolution of singularities, and showing $\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)$ is reflexive [GK14a].
 By Flenner’s criterion, this is novel exactly when $\mathrm{codim}_{X}(\Sigma)=2$.
@@
 […]
-Recalling that general hyperplane sections preserve the du Bois property, an inductive hypothesis would imply that the left and right vertical morphisms are isomorphisms — this is exactly the approach used in [GK14] for $p=1$, but they need to use the negativity lemma [GKK10, Proposition 7.5], which seems particularly special for $p=1$.
+Recalling that general hyperplane sections preserve the Du Bois property, an inductive hypothesis would imply that the left and right vertical morphisms are isomorphisms — this is exactly the approach used in [GK14a] for $p=1$, but they need to use the negativity lemma [GKK10, Proposition 7.5], which seems particularly special for $p=1$.
 Induction seems insufficient in proving $\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is an isomorphism, and the reflexivity of the sheaf $(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p}(\log E|_{H})$ is not sufficient to prove the reflexivity of $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$.
@@
 We will often apply this to a singular variety $X$ of dimension $n$ by considering a local embedding $X\subset Y$ into a smooth complex manifold $Y$.
-To highlight the idea, let $X$ be a normal variety of dimension $n$ with at worst isolated du Bois singularities, which implies that the cohomology sheaves $R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$ for $i>0$ for any log-resolution of singularities $\pi:\widetilde{X}\to X$.
-By Grauert-Rimenschneider vanishing, there is an induced quasi-isomorphism
-(1.6) $\pi_{*}\omega_{\widetilde{X}}(E)\cong\mathbf{R}\pi_{*}\omega_{\widetilde{X}}(E),$
-and Grothendieck duality immediately implies that
-$R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(\pi_{*}\omega_{\widetilde{X}}(E),\omega_{X}^{\bullet})[-n]\cong R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$
-for $j>0$.
-By Proposition 1.3, $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive.
-We emphasize that this vanishing is much stronger than logarithmic extension and is only equivalent if $X$ is Cohen-Macaulay.
-Theorem 1.1 follows more generally from the following lemma:
-Lemma 1.4.
-If $X$ is a normal complex algebraic variety with a log-resolution of singularities $\pi:\widetilde{X}\to X$, then $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{0})\cong\pi_{*}\omega_{\widetilde{X}}(E)$, where $\mathbb{D}_{X}$ is the Grothendieck duality functor.
-For holomorphic extension, the problem is more subtle. We need to extend Proposition 1.3 to the level of complexes:
+There is also a derived version of this criterion discussed in [KS21].
+We remark that this statement is stronger than holomorphic extension:
 Proposition 1.4.
@@
 for every $k\in\mathbb{Z}$, then the sections of $\mathscr{H}^{0}K$ extend uniquely across $A$.
+To highlight the idea, let $X$ be a normal variety of dimension $n$ with at worst isolated Du Bois singularities, which implies that the cohomology sheaves $R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$ for $i>0$ for any log-resolution of singularities $\pi:\widetilde{X}\to X$.
+We use Proposition 1.4 to test holomorphic extension on
+(1.6) $\pi_{*}\omega_{\widetilde{X}}(E)=R^{0}\pi_{*}\omega_{\widetilde{X}}(E).$
+By Grothendieck duality,
+$R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(\mathbf{R}\pi_{*}\omega_{\widetilde{X}}(E),\omega_{X}^{\bullet})[-n]\cong R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$
+for $j>0$.
+Thus, $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive.
+We emphasize that this vanishing is much stronger than logarithmic extension and is only equivalent if $X$ is Cohen-Macaulay.
+Theorem 1.1 follows more generally from the following lemma:
+Lemma 1.5.
+If $X$ is a normal complex algebraic variety with a log-resolution of singularities $\pi:\widetilde{X}\to X$, then $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{0})\cong\pi_{*}\omega_{\widetilde{X}}(E)$, where $\mathbb{D}_{X}$ is the Grothendieck duality functor.
+For holomorphic extension, the problem is more subtle.
 The naive approach is to consider the support of the sheaves $R^{j}\pi_{*}\Omega_{\widetilde{X}}^{\dim X-p}$ for $p=\mathrm{codim}_{X}(\Sigma)$, but this seems hopeless: for instance, $R^{1}\pi_{*}\Omega_{\widetilde{X}}^{1}$ is non-zero for ADE surface singularities — these singularities satisfy holomorphic extension in all degrees, either classically or by [KS21, Corollary 1.8].
@@
 Let $X$ be a normal variety with singular locus $\Sigma$. The inclusion morphism (1.4) is an isomorphism for $p=\mathrm{codim}_{X}(\Sigma)-1$.
-1.6. Holomorphic Extension for $k$-du Bois Singularities
+1.6. Holomorphic Extension for $m$-Du Bois Singularities
 Du Bois singularities are of fundamental interest to algebraic geometers as a general class of singularities for which deformation theory and Hodge theory can be studied.
-It is interesting to consider $k$-du Bois singularites, the class of singularities which restricts the higher graded pieces of the du Bois complex
-We say that a normal complex algebraic variety $X$ has weakly (or pre-) $k$-du Bois singularities if the cohomology sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}=0$ for $j>0$ and $0\leq p\leq k$, where $\underline{\Omega}_{X}^{p}:=\mathrm{gr}_{F}^{p}\underline{\Omega}_{X}^{\bullet}[p]$.
-Alternatively, $X$ has weakly $k$-du Bois singularities if the natural map
+It is interesting to consider $m$-Du Bois singularities, the class of singularities which restricts the higher graded pieces of the Du Bois complex.
+We say that a normal complex algebraic variety $X$ has weakly (or pre-) $m$-Du Bois singularities if the cohomology sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}=0$ for $j>0$ and $0\leq p\leq m$, where $\underline{\Omega}_{X}^{p}:=\mathrm{gr}^{F}_{p}\underline{\Omega}_{X}^{\bullet}[p]$.
+Alternatively, $X$ has weakly $m$-Du Bois singularities if the natural map
 $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$
-is a quasi-isomorphism. This notion has been studied in [SVV23] and is a generalization of the notion of $k$-du Bois singularities appearing in [MOPW21], [JKSY21], and [MP22].
-An interesting question is then to what extent Theorem 1.2 can be improved for weakly $k$-du Bois singularities.
-For hypersurfaces, the answer is known: a variety $X$ with hypersurface singularities satisfies holomorphic extension in degree $\dim X$ if and only if $X$ is 1-du Bois, which means that $X$ is du Bois and the natural map
+is a quasi-isomorphism. This notion has been studied in [SVV23] and is a generalization of the notion of $m$-Du Bois singularities appearing in [MOPW21], [JKSY21], and [MP22].
+An interesting question is then to what extent Theorem 1.2 can be improved for weakly $m$-Du Bois singularities.
+For hypersurfaces, the answer is known: a variety $X$ with hypersurface singularities satisfies holomorphic extension in degree $\dim X$ if $X$ is 1-Du Bois, which means that $X$ is Du Bois and the natural map
 $\Omega_{X}^{1}\xrightarrow{\sim}\underline{\Omega}_{X}^{1}$
 is a quasi-isomorphism, where $\Omega_{X}^{1}$ is the sheaf of Kähler 1-forms.
-This seems to be rather special for hypersurface (or more generally complete intersection) singularities, as we can write down a Cohen-Macaulay variety with weakly $k$-du Bois singularities and $k>0$ in any dimension for which Theorem 1.2 is still optimal (see Example 6.1).
+This seems to be rather special for hypersurface (or more generally complete intersection) singularities, as we can write down a Cohen-Macaulay variety with weakly $m$-Du Bois singularities and $k>0$ in any dimension for which Theorem 1.2 is still optimal (see Example 6.1).
 The algebraic properties of the sheaf $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ appear to be independent of the vanishing — or non-vanishing — of the sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}$.
@@
 Theorem 1.8.
-Let $X$ be a normal variety with weakly $k$-du Bois singularities, and let $\pi:\widetilde{X}\to X$ be a resolution of singularities.
+Let $X$ be a normal variety with weakly $m$-Du Bois singularities, and let $\pi:\widetilde{X}\to X$ be a resolution of singularities.
 If $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then $\pi_{*}\Omega_{\widetilde{X}}^{k+1}$ is reflexive.
@@
 Let $X$ be a normal complex algebraic variety.
-Following [MOPW21] and [JKSY21], we say that $X$ has $k$-du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $p$, where $\Omega_{X}^{p}$ is the sheaf of Kähler $p$-forms.
+Following [MOPW21] and [JKSY21], we say that $X$ has $m$-Du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $p$, where $\Omega_{X}^{p}$ is the sheaf of Kähler $p$-forms.
 As a generalization of Du Bois singularities, this is very natural, as it comes from a morphism of complexes $\Omega_{X}^{\bullet}\to\underline{\Omega}_{X}^{\bullet}$.
 However, this definition is rather restrictive: in the case of hypersurface singularities, this forces $\Omega_{X}^{p}$ to be reflexive for every $0\leq p\leq k$ and restricts the codimension of the singularities.
-Instead, one can consider what happens when we restrict the duals of the higher du Bois complexes. We say that has weakly $k$-rational singularities if
+Instead, one can consider what happens when we restrict the duals of the higher Du Bois complexes. We say that has weakly $m$-rational singularities if
 $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$
@@
 Note that there is a quasi-isomorphism $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n})\cong\mathbf{R}\pi_{*}\mathscr{O}_{\widetilde{X}}$, whence 00-rational is the same as rational singularities for normal complex varieties.
-This is a generalization of the $k$-rational definition established for complete intersections, which requires the natural morphism $\Omega_{X}^{p}\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ to be an isomorphism for $0\leq p\leq k$.
-As was observed in [MOPW21], the $k$-du Bois definition restricts the codimension of the singularities of a hypersurface for large $k$, and $X$ is smooth if and only if $k>\frac{1}{2}\dim X$.
-We generalize this result for $k$-rational singularities.
-We remark that this appears to be the correct generalization, as $k$-du Bois singularities are $(k-1)$-rational for hypersurfaces by [FL22b].
-Theorem 1.9.
-If $X$ is a normal complex algebraic variety with $k$-rational and Gorenstein singularities, then $k\leq\frac{1}{2}\mathrm{codim}_{X}(\Sigma)$.
-1.8. Hodge Theory of Weakly $k$-Rational Singularities
+This is a generalization of the $m$-rational definition established for complete intersections, which requires the natural morphism $\Omega_{X}^{p}\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ to be an isomorphism for $0\leq p\leq k$.
+The $m$-rational property is very strong, as it eventually implies $\Omega_{X}^{1}$ is maximal Cohen-Macaulay.
+The MCM property has been extensively studied in the literature, and the MCM property for the module of Kähler differentials has been looked at in the case of hypersurfaces and when $\mathrm{pd}~{}\Omega_{X,x}^{p}<\infty$.
+We give one new result using the theory of Hodge modules.
+Proposition 1.9.
+If $X$ is an $n$-fold Gorenstein variety with at worst quotient singularities, then $\Omega_{X}^{p}$ is reflexive for some $1\leq p\leq n$ if and only if $X$ is smooth.
+1.8. Hodge Theory of Weakly $m$-rational Singularities
 As we mentioned above, the Du Bois complex is a Hodge-theoretic object arising from Deligne’s mixed Hodge theory.
@@
 $E_{1}^{p,q}=\mathbb{H}^{q}(X,\underline{\Omega}_{X}^{p})\Rightarrow H^{p+q}(X,\mathbb{C}).$
-If $X$ is weakly $k$-du Bois, we get a decomposition
-$H^{k}(X,\mathbb{C})=\bigoplus_{p+q=k}H^{q}(X,\mathscr{H}^{0}\underline{\Omega}_{X}^{p}).$
+If $X$ is weakly $m$-Du Bois, we get a decomposition
+$H^{k}(X,\mathbb{C})\cong\bigoplus_{p+q=k}H^{q}(X,\mathscr{H}^{0}\underline{\Omega}_{X}^{p}).$
 In general, restricting the Du Bois complex does not affect the weight filtration on the cohomology, as $\underline{\Omega}_{X}^{\bullet}$ does not usually admit the structure of a Hodge module.
-To get pure Hodge modules, we consider again $k$-rational definition.
+To get pure Hodge modules, we consider again $m$-rational definition.
 Theorem 1.10.
-If $X$ is a normal complex algebraic variety with weakly $k$-rational singularities, then the Hodge filtration on $H^{k}(X,\mathbb{Q})$ induces a pure Hodge structure.
+If $X$ is a normal and proper complex algebraic variety with weakly $m$-rational singularities, then the Hodge filtration on $H^{k}(X,\mathbb{Q})$ induces a pure Hodge structure.
 1.9. Acknowledgements
@@
 2.1. Notation
-Throughout, we let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$.
-We will use $\pi:\tilde{X}\to X$ to denote a strong and projective resolution of singularities.
+Throughout, we let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$. We will use $\pi:\tilde{X}\to X$ to denote a projective resolution of singularities.
 If $\pi:\tilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, we let $\Omega_{\tilde{X}}^{p}(\log E)$ be the sheaf of log $p$-forms, and we let
@@
 Let $X$ be a complex algebraic variety.
-The du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$ is an object in the derived category of filtered complexes generalizing the holomorphic de Rham complex for algebraic varieties over $\mathbb{C}$.
-We denote its graded pieces by $\underline{\Omega}_{X}^{p}:=\mathrm{gr}_{F}^{p}\underline{\Omega}_{X}^{\bullet}[p]$.
+The Du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$ is an object in the derived category of filtered complexes of constructible sheaves $\mathscr{D}^{b}_{\mathrm{filt}}(X)$, generalizing the holomorphic de Rham complex for algebraic varieties over $\mathbb{C}$.
+We denote its graded pieces by $\underline{\Omega}_{X}^{p}:=\mathrm{gr}^{F}_{p}\underline{\Omega}_{X}^{\bullet}[p]$, which are defined in the bounded derived category of coherent sheaves.
 Studied by Du Bois [DB81] from Deligne’s construction of the mixed Hodge structure [Del74], the Du Bois complex is constructed by simplicial or cubical hyperresolutions of $X$.
@@
   6. (vi)
-    [Kov11, Proposition 2.6], [SVV23, Lemma 3.2] If $H\subset X$ is a general member of a basepoint free linear system, then $\underline{\Omega}_{H}^{\bullet}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{\bullet}\otimes\mathscr{O}_{H}$.
+    [Kov11, Proposition 2.6], [SVV23, Lemma 3.2] If $H\subset X$ is a general member of a basepoint free linear system, then there is an exact triangle
+    $\underline{\Omega}_{H}^{p-1}\otimes_{\mathscr{O}_{H}}^{\mathbb{L}}\mathscr{O}_{H}(-H)\to\underline{\Omega}_{X}^{p}\otimes_{\mathscr{O}_{X}}^{\mathbb{L}}\mathscr{O}_{H}\to\underline{\Omega}_{H}^{p}\xrightarrow{+1}.$
+    In particular, $\underline{\Omega}_{X}^{0}\otimes\mathscr{O}_{H}\cong_{\mathrm{qis}}\underline{\Omega}_{H}^{0}$.
   7. (vii)
@@
     $\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{\Sigma}^{p}\oplus\mathbf{R}\pi_{*}\Omega_{\tilde{X}}^{p}\to\mathbf{R}\pi_{*}\underline{\Omega}_{E}^{p}\xrightarrow{+1}$
-    where $\pi:\tilde{X}\to X$ is a morphism of complex schemes of finite type with singular locus $\Sigma\subset X$ and $E=\pi^{-1}(\Sigma)$.
-In particular, $\underline{\Omega}_{X}^{n}\cong_{\mathrm{qis}}\mathbf{R}\pi_{*}\omega_{\tilde{X}}\cong_{\mathrm{qis}}\pi_{*}\omega_{\tilde{X}}$ if $\pi:\tilde{X}\to X$ is a resolution of singularities.
+    where $\pi:\tilde{X}\to X$ is a resolution of singularities, $\Sigma\subset X$ is the singular locus, and $E=\pi^{-1}(\Sigma)$.
+In particular, $\underline{\Omega}_{X}^{n}\cong_{\mathrm{qis}}\mathbf{R}\pi_{*}\omega_{\tilde{X}}\cong_{\mathrm{qis}}\pi_{*}\omega_{\tilde{X}}$.
   8. (viii)
@@
     which is independent of the choice of $\pi$.
-2.3. du Bois, $k$-du Bois, and $k$-rational Singularities
+2.3. Du Bois, $m$-Du Bois, and $m$-rational Singularities
 Let $X$ be a complex algebraic variety. We review classes of singularities which are associated to the complexes $\underline{\Omega}_{X}^{p}$.
@@
   2. (ii)
-    We say that $X$ has $k$-du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq k$.
+    We say that $X$ has $m$-Du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq m$.
   3. (iii)
-    We say that $X$ has $k$-rational singularities if the natural map $\Omega_{X}^{p}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ is a quasi-isomorphism for each $0\leq p\leq k$.
+    We say that $X$ has $m$-rational singularities if the natural map $\Omega_{X}^{p}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ is a quasi-isomorphism for each $0\leq p\leq m$.
   4. (iv)
-    We say that $X$ has weakly $k$-du Bois singularities, or pre $k$-du Bois singularities, if the natural map $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq k$.
+    We say that $X$ has weakly $m$-Du Bois singularities if the natural map $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $0\leq p\leq m$.
   5. (v)
-    We say that $X$ has weakly $k$-rational singularities, or pre $k$-rational singularities, if the natural map $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ for each $0\leq p\leq k$.
-Here are some general properties of (weakly) $k$-du Bois (resp. (weakly) $k$-rational singularities).
-  • By Theorem 2.1(vi), if $H$ is a general member of a basepoint linear system of a variety $X$ with (weakly) $k$-du Bois singularities, then $H$ also has (weakly) $k$-du Bois singularities (resp.
-(weakly) $k$-rational singularities).
+    We say that $X$ has weakly $m$-rational singularities if the natural map $\mathscr{H}^{0}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ for each $0\leq p\leq m$.
+Here are some general properties of (weakly) $m$-Du Bois (resp. (weakly) $m$-rational singularities).
+  • If $H$ is a general member of a basepoint linear system of a variety $X$ with (weakly) $m$-Du Bois singularities, then $H$ also has (weakly) $m$-Du Bois singularities (resp.
+(weakly) $m$-rational singularities) [SVV23, Theorem A, Corollary 3.3].
   • Rational and log-canonical singularities are Du Bois, see [Kov99, Corollary 2.6], [Sai00, 5.4. Theorem], and [KK10, Theorem 1.4].
-  • If $X$ is a normal variety and $f:Y\to X$ is a finite dominant map from a variety $Y$ with rational and weakly $k$-du Bois singularities, then $X$ also has rational [Kov00, Theorem 1] and weakly $k$-du Bois (resp.
-weakly $k$-rational) singularities [SVV23, Proposition 4.2].
-  • If $X$ has simple normal crossing singularities, then $X$ is weakly $k$-du Bois since
+  • If $X$ is a normal variety and $f:Y\to X$ is a finite dominant map from a variety $Y$ with rational and weakly $m$-Du Bois singularities, then $X$ also has rational [Kov00, Theorem 1] and weakly $m$-Du Bois (resp.
+weakly $m$-rational) singularities [SVV23, Proposition 4.2].
+  • If $X$ has simple normal crossing singularities, then $X$ is weakly $m$-Du Bois since
     $\underline{\Omega}_{X}^{p}\cong\Omega_{X}^{p}/\mathrm{tor}$
@@
   • If $X$ has rational singularities, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\cong\Omega_{X}^{[p]}$ for each $p$, which follows from [KS21] and [HJ14].
-In particular, a complex algebraic variety $X$ with rational singularities is weakly $k$-du Bois (resp.
-$k$-rational) if and only if $\Omega_{X}^{[p]}\to\underline{\Omega}_{X}^{p}$ (resp.
+In particular, a complex algebraic variety $X$ with rational singularities is weakly $m$-Du Bois (resp.
+$m$-rational) if and only if $\Omega_{X}^{[p]}\to\underline{\Omega}_{X}^{p}$ (resp.
 $\Omega_{X}^{[p]}\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$) is a quasi-isomorphism for each $0\leq p\leq k$.
-There is a nice characterization due to Schwede which wish to mention.
+There is a nice characterization due to Schwede.
 For $X$ a reduced and separated complex scheme of finite type, there exists a (local) embedding $\iota:X\to Y$ of $X$ into a smooth scheme $Y$.
@@
 Let $Y$ be a smooth complex manifold of dimension $d$ (for example, let $Y$ be a local embedding of a complex algebraic variety $X$ of codimension $c=d-n$).
-A pure Hodge module on $Y$ is an object $M=(\mathcal{M},F_{\bullet},\mathrm{rat}~{}M)$ consisting of:
+A pure Hodge module on $Y$ is an object $M=(\mathcal{M},F^{\bullet},\mathrm{rat}~{}M)$ consisting of:
   1. (i)
@@
     $F_{p}\mathcal{M}\cdot F_{q}\mathscr{D}_{Y}\subset F_{p+q}\mathcal{M},$
-    and $\mathrm{gr}_{F}^{\bullet}$ is coherent $\mathrm{gr}_{F}^{\bullet}\mathscr{D}_{Y}$.
+    and $\mathrm{gr}^{F}_{\bullet}\mathcal{M}$ is coherent over $\mathrm{gr}^{F}_{\bullet}\mathscr{D}_{Y}$.
   3. (iii)
@@
 Because of this, we can define the category $\mathrm{HM}(Y,w)$ of polarized Hodge modules of weight $w$ by inducing a polarization from the data $M_{Y_{i}}$ of polarized variations of pure Hodge structures.
-A mixed Hodge module on $Y$ is an object $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)$ consisting of a $\mathscr{D}_{Y}$-module structure $\mathcal{M}$, a perverse structure $\mathrm{rat}~{}M$, a decreasing good filtration $F_{\bullet}$ on $\mathcal{M}$, and an increasing weight filtration $W_{\bullet}$ on these structures such that the graded pieces
-$\mathrm{gr}_{k}^{W}M=(\mathrm{gr}_{k}^{W}\mathcal{M},\mathrm{rat}~{}\mathrm{gr}_{k}^{W}M,F_{\bullet})$
+A mixed Hodge module on $Y$ is an object $M=(\mathcal{M},F^{\bullet},W_{\bullet},\mathrm{rat}~{}M)$ consisting of a $\mathscr{D}_{Y}$-module structure $\mathcal{M}$, a perverse structure $\mathrm{rat}~{}M$, a decreasing good filtration $F^{\bullet}$ on $\mathcal{M}$, and an increasing weight filtration $W_{\bullet}$ on these structures such that the graded pieces
+$\mathrm{gr}_{W}^{k}M=(\mathrm{gr}_{W}^{k}\mathcal{M},\mathrm{rat}~{}\mathrm{gr}_{W}^{k}M,F^{\bullet})$
 are pure Hodge modules.
-We call $M$ a graded-polarizable mixed Hodge module if further $\mathrm{gr}_{k}^{W}M$ is polarizable for each $k$.
+We call $M$ a graded-polarizable mixed Hodge module if further $\mathrm{gr}_{W}^{k}M$ is polarizable for each $k$.
 We refer to the category of graded-polarizable mixed Hodge modules as $\mathrm{MHM}(Y)$.
 3.1.2. The Dual Mixed Hodge Module
-Let $Y$ be a complex manifold of dimension $d$, and fix a mixed Hodge module $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$.
-We define the dual Hodge module $\mathbb{D}_{Y}(M)$ to be the mixed Hodge module satisfying
-(3.2) $\mathbb{D}_{Y}(\mathrm{gr}_{k}^{W}M)=\mathrm{gr}_{-k}^{W}\mathbb{D}_{Y}(M)$
-for every $k$. The underlying perverse sheaf $\mathrm{rat}~{}\mathbb{D}(M)$ is the Verdier dual of $\mathrm{rat}~{}M$, and the underlying $\mathscr{D}_{Y}$-module is the holonomic dual
-$\mathbb{D}_{Y}(\mathcal{M})=R^{d}\mathrm{Hom}_{\mathscr{D}_{Y}}(\omega_{Y}\otimes_{\mathscr{O}_{Y}}\mathcal{M},\mathscr{D}_{Y}),$
-which is compatible with the filtration $F_{\bullet}$ of item (ii) of §3.1.1, by [Sai88, Lemma 5.1.13].
-In particular, if $M$ is a pure Hodge module of weight $w$, then $\mathbb{D}_{Y}(M)$ is a pure Hodge module of weight $-w$.
-In this case, $\mathbb{D}_{Y}(M)\cong M(w)$, where $M(w)$ is the Tate-twist of $M$ in degree $w$, obtained by twisting the perverse structure and Hodge filtration in the usual way.
-The compatibility of the Hodge filtration with $\mathbb{D}_{Y}$ is due to Saito [Sai88, 2.4.3]. Specifically, there is an isomorphism
-(3.3) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{M})),$
-where again $\omega_{Y}^{\bullet}=\omega_{Y}[d]$. If $M\in\mathrm{HM}(Y,w)$, the isomorphism $\mathbb{D}_{Y}(M)\cong M(w)$ reduces (3.3) to
+Let $Y$ be a complex manifold of dimension $d$, and fix a mixed Hodge module $M=(\mathcal{M},F^{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$.
+There is a mixed Hodge module $\mathbf{D}_{Y}(M)$, known as the dual Hodge module, satisfying
+(3.2) $\mathbf{D}_{Y}(\mathrm{gr}_{W}^{k}M)=\mathrm{gr}_{-k}^{W}\mathbf{D}_{Y}(M)$
+for every $k$. The underlying perverse sheaf $\mathrm{rat}~{}\mathbf{D}_{Y}(M)$ is the Verdier dual of $\mathrm{rat}~{}M$, and the underlying $\mathscr{D}_{Y}$-module is the holonomic dual
+$\mathbf{D}_{Y}(\mathcal{M})=R^{d}\mathrm{Hom}_{\mathscr{D}_{Y}}(\omega_{Y}\otimes_{\mathscr{O}_{Y}}\mathcal{M},\mathscr{D}_{Y}),$
+which is compatible with the filtration $F^{\bullet}$ of item (ii) of §3.1.1, by [Sai88, Lemma 5.1.13].
+In particular, if $M$ is a pure Hodge module of weight $w$, then $\mathbf{D}_{Y}(M)$ is a pure Hodge module of weight $-w$.
+In this case, $\mathbf{D}_{Y}(M)\cong M(w)$, where $M(w)$ is the Tate-twist of $M$ in degree $w$, obtained by twisting the perverse structure and Hodge filtration in the usual way.
+The compatibility of the Hodge filtration with $\mathbf{D}_{Y}$ is due to Saito [Sai88, 2.4.3]. Specifically, there is an isomorphism
+(3.3) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{M})),$
+where again $\omega_{Y}^{\bullet}=\omega_{Y}[d]$. If $M\in\mathrm{HM}(Y,w)$, the isomorphism $\mathbf{D}_{Y}(M)\cong M(w)$ reduces (3.3) to
 (3.4) $\mathbf{R}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathcal{M}),\omega_{Y}^{\bullet})\cong\mathrm{gr}_{-p-w}^{F}\mathrm{DR}(\mathcal{M}).$
 3.1.3. The Restricted Hodge Module
-Now let $Y=\mathbb{C}^{d}$ and $M=(\mathcal{M},F_{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$.
+Now let $Y=\mathbb{C}^{d}$ and $M=(\mathcal{M},F^{\bullet},W_{\bullet},\mathrm{rat}~{}M)\in\mathrm{MHM}(Y)$.
 For a generic hyperplane section $H$ of $Y$, $H$ intersects any Whitney strata adapted to the perverse sheaf $\mathrm{DR}(\mathcal{M})$ transversely.
@@
 $\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}\mathcal{M}$
-with filtration $F_{\bullet}\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}F_{\bullet}\mathcal{M}$, see [Sch16, Lemma 9.5].
+with filtration $F^{\bullet}\mathcal{M}_{H}=\mathscr{O}_{H}\otimes_{\iota_{H}^{-1}\mathscr{O}_{Y}}\iota_{H}^{-1}F^{\bullet}\mathcal{M}$, see [Sch16, Lemma 9.5].
 The de Rham complex of $\mathcal{M}_{H}$ is
@@
 The intersection Hodge module $\mathrm{IC}_{X}=(\mathcal{IC}_{X},F,IC_{X})\in\mathrm{HM}(Y,n)$ is the unique Hodge module with support $X$ determined by the variation of Hodge structures $\mathbb{Q}_{U}[n]$, where $U=X_{\mathrm{reg}}$ is the regular locus of $X$.
-We remark that the underlying perverse sheaf $IC_{X}$ of $\mathbb{Q}$-constructible sheaves computes the intersection cohomology groups $IH^{\bullet}(X,\mathbb{Q}):=\mathbb{H}^{\bullet-n}(X,IC_{X})$ of $X$.
+We remark that the underlying perverse sheaf $IC_{X}$ computes the intersection cohomology $IH^{\bullet}(X,\mathbb{Q}):=\mathbb{H}^{\bullet-n}(X,IC_{X})$ of $X$.
 With this in mind, we get a Hodge-theoretic interpretation of the decomposition theorem.
 Let $\pi:\widetilde{X}\to X$ be a resolution of singularities, and consider the induced morphism $f=\iota\circ\pi:\widetilde{X}\to Y$.
-By Saito’s direct image theorem [Sai88, §5.3], there are pure Hodge modules $M_{l}=(\mathcal{M}_{l},F_{\bullet},\mathrm{rat}~{}M_{l})$ supported on the singularities of $X$ and a decomposition
+By Saito’s direct image theorem [Sai88, §5.3], there are pure Hodge modules $M_{l}=(\mathcal{M}_{l},F^{\bullet},\mathrm{rat}~{}M_{l})$ supported on the singularities of $X$ and a decomposition
 (3.6) $\mathbf{R}f_{*}\Omega_{\widetilde{X}}^{p}[n-p]\cong_{\mathrm{qis}}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})\oplus\bigoplus_{l\in\mathbb{Z}}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{M}_{l})[-l],$
@@
 If $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with log-exceptional divisor $E$, recall from §2.1 that we assume an isomorphism $U:=X_{\mathrm{reg}}\cong\widetilde{X}\setminus E$.
-The sheaf $\mathbb{Q}_{\widetilde{X}\setminus E}$ defines a variation of pure Hodge structures; as a perverse sheaf, we can consider the pushforward $\mathbf{R}j_{*}\mathbb{Q}_{\widetilde{X}\setminus E}$, where $j:\tilde{X}\setminus E\hookrightarrow\widetilde{X}$ is the inclusion.
+The sheaf $\mathbb{Q}_{\widetilde{X}\setminus E}[n]$ defines a variation of pure Hodge structures; as a perverse sheaf, we can consider the pushforward $\mathbf{R}j_{*}\mathbb{Q}_{\widetilde{X}\setminus E}$, where $j:\tilde{X}\setminus E\hookrightarrow\widetilde{X}$ is the inclusion.
 By [Sai90, Thm.
@@
 We refer to the underlying $\mathscr{D}_{Y}$-module as $\mathcal{N}_{0}$.
-By [KS21, Proposition 9.5], we get the following important relationship between logarithmic forms and the logarithmic Hodge module $N_{0}=(\mathcal{N}_{0},F_{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$:
+By [KS21, Proposition 9.5], we get the following important relationship between logarithmic forms and the logarithmic Hodge module $N_{0}=(\mathcal{N}_{0},F^{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$:
 (3.8) $f_{*}\Omega_{\widetilde{X}}^{p}(\log E)\cong\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0}).$
@@
     Logarithmic extension holds in degree $p\geq 0$ if
-    (3.10) $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0}))\leq n-j-p-2$
+    (3.10) $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0}))\leq n-j-p-2$
     for every $j\geq 0$, where $\mathcal{N}_{0}$ is the $\mathscr{D}_{Y}$-module underlying the logarithmic Hodge module $N_{0}$.
@@
 $E_{2}^{p,q}=R^{p}\mathscr{H}om_{\mathscr{O}_{Y}}(\mathscr{H}^{-q}\underline{\Omega}_{\Sigma}^{0},\omega_{Y}^{\bullet})\Rightarrow R^{p+q}\mathscr{H}om_{\mathscr{O}_{Y}}(\underline{\Omega}_{\Sigma}^{0},\omega_{Y}^{\bullet}).$
-Note that $\mathrm{Supp}~{}E_{2}^{p,q}\neq 0$ if $j=-\dim\Sigma,...,0$, whence $\dim\mathrm{Supp}~{}E_{2}^{p,q}=0$ if $p<\dim Z-q$ by Theorem 2.1(iv).
+Note that $\mathrm{Supp}~{}E_{2}^{p,q}\neq 0$ if $j=-\dim\Sigma,...,0$, whence $\dim\mathrm{Supp}~{}E_{2}^{p,q}=0$ if $p<\dim\Sigma-q$ by Theorem 2.1(iv).
 Thus $E_{\infty}^{p,q}=0$ in this range, which gives the desired vanishing on $X$.
@@
 Proof of Theorem 1.1.
-By assumption, $\underline{\Omega}_{X}^{0}$ has no higher cohomology, and $\dim\mathrm{Supp}~{}\mathscr{O}_{X}<n-2$ since $X$ is normal.
+By assumption, $\underline{\Omega}_{X}^{0}$ has no higher cohomology, and $\dim\Sigma\leq n-2$ since $X$ is normal.
 Therefore logarithmic extension holds in degree $n$ by Lemma 4.1 and Proposition 1.4.
@@
 If $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive, we get a family of inequalities
-$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))\leq-(j+p+2)$
+$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))\leq-(j+p+2)$
 for $p+j\geq-n+1$, see [KS21, Proposition 9.10]. This leads to an important vanishing as a consequence of Theorem 1.1.
@@
 Let $X$ be a normal complex analytic variety of dimension $n$ with at worst Du Bois singularities. Then
-$\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))=0,$
+$\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))=0,$
 where $X\subset Y$ is a (locally) closed embedding into a smooth manifold $Y$ and $\mathcal{N}_{0}$ is the $\mathscr{D}_{Y}$-module underlying the logarithmic mixed Hodge module.
@@
 Proposition 5.2.
-Let $X$ be a normal variety of dimension $n$ with at worst du Bois singularities, and let $\mathrm{IC}_{X}=(\mathcal{I}C_{X},F_{\bullet},IC_{X})$ be the intersection Hodge module. Then
+Let $X$ be a normal variety of dimension $n$ with at worst Du Bois singularities, and let $\mathrm{IC}_{X}=(\mathcal{I}C_{X},F^{\bullet},IC_{X})$ be the intersection Hodge module. Then
 $\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathcal{IC}_{X})=0.$
@@
 Let $X\subset Y$ be a (locally) closed embedding into a smooth complex manifold $Y$.
-Let $N_{0}=(\mathcal{N}_{0},F_{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$ be the logarithmic mixed Hodge module.
-By construction, $W_{n}\mathbb{D}_{Y}(\mathcal{N}_{0})=\mathbb{D}_{Y}(W_{n}\mathcal{N}_{0})$.
+Let $N_{0}=(\mathcal{N}_{0},F^{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$ be the logarithmic mixed Hodge module.
+By construction, $\mathrm{gr}_{-n}\mathbf{D}_{Y}(\mathcal{N}_{0})=\mathbf{D}_{Y}(W_{n}\mathcal{N}_{0})$.
 Since $\mathrm{gr}_{-1}^{F}\mathrm{DR}(-)$ is an exact functor, there is an exact sequence
-$\begin{split}\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))&\to\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}(\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))\\ &\to\mathscr{H}^{1}\mathrm{gr}_{-1}^{F}\mathrm{DR}(W_{n-1}\mathbb{D}(\mathcal{N}_{0})(-n)).\end{split}$
-The last term must be zero for degree reasons, and so by Corollary 5.1 we get the additonal vanishing $\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))=0$.
-By (3.3), we have an isomorphism $\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n)\cong\mathrm{gr}_{n}^{W}\mathcal{N}_{0}$.
+$\begin{split}\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))&\to\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}(\mathrm{gr}_{n}^{W}\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))\\ &\to\mathscr{H}^{1}\mathrm{gr}_{-1}^{F}\mathrm{DR}(W_{n-1}\mathbb{D}(\mathcal{N}_{0})(-n)).\end{split}$
+The last term must be zero for degree reasons, and so by Corollary 5.1 we get the additonal vanishing $\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathrm{gr}_{n}^{W}\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))=0$.
+By (3.3), we have an isomorphism $\mathrm{gr}_{n}^{W}\mathbf{D}_{Y}(\mathcal{N}_{0})(-n)\cong\mathrm{gr}_{n}^{W}\mathcal{N}_{0}$.
 The claim follows from the isomorphism
@@
 Let $\mathrm{IC}_{X}$ and $N_{0}$ be the intersection and logarithmic Hodge modules of the normal variety $X$.
-We note there is a canonical morphism of (mixed) Hodge modules $M_{X}\to N_{0}$ obtained by the isomorphism $W_{n}N_{0}\cong\mathrm{IC}_{X}$ [KS21, Proof of (9.8.1)].
+We note there is a canonical morphism of (mixed) Hodge modules $\mathcal{IC}_{X}\to N_{0}$ obtained by the isomorphism $W_{n}N_{0}\cong\mathrm{IC}_{X}$ [KS21, Proof of (9.8.1)].
 The problem is local, so let $X\subset Y$ be a closed embedding of $X$ into a ball $Y\cong\mathbb{C}^{n+c}$, and let $H\subset Y$ be a generic hyperplane section.
 For each $p$, there is a commutative diagram of exact sequences
-[…]
+00{0}00{0}${N_{H|Y}^{*}\otimes_{\mathscr{O}_{H}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})}$${N_{H|Y}^{*}\otimes_{\mathscr{O}_{H}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{N}_{0}|_{H})}$${\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})}$${\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0}|_{H})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{N}_{0}|_{H})}$
 coming from the exact triangle associated to the restricted Hodge module sequence (3.5).
@@
 For this $p$, we also have
-$\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})=\mathscr{H}^{-\dim\Sigma}\mathrm{gr}_{-((n-1)-\dim\Sigma)}\mathrm{DR}(\mathcal{IC}_{X\cap H}).$
-For dimension reasons, this term vanishes. Therefore the morphism
-$\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$
-is an isomorphism by the 4-lemma. The theorem follows since $H$ was chosen to be general. ∎
+$\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})=\mathscr{H}^{-\dim\Sigma}\mathrm{gr}^{F}_{-((n-1)-\dim\Sigma)}\mathrm{DR}(\mathcal{IC}_{X\cap H}).$
+For dimension reasons, this term vanishes, and we see the map
+$\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}\to(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p}$
+is surjective.
+Next, we follow an argument in [GK14a].
+Let $E=\sum E_{i}$ be the exceptional divisor of $\pi$, and let $\alpha$ be section in $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{p}(\log E))$.
+By definition, there are integers $m_{i}\in\left\{0,1\right\}$ such that $\gamma$ is a section in $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{p}(\sum m_{i}E_{i}))$.
+Consider the diagram
+[…]
+We assume again the assumption is true in dimension $\dim X-1$.
+In particular, the image $\beta(\alpha(\gamma))$ factors through $H^{0}(\widetilde{H},\Omega_{\widetilde{H}}^{p})$.
+Supposing by contradiction that the $m_{i}$ are not all 0, this implies the existence of a nonzero section of $H^{0}(\widetilde{H},N_{\widetilde{H}|\widetilde{X}}^{*}\otimes\Omega_{\widetilde{X}}^{p-1}\otimes(\sum m_{i}E_{i}))$.
+Since the pullback of $N_{H|X}$ is just $N_{\widetilde{H}|\widetilde{X}}$, this implies $H^{0}(\widetilde{H},\Omega_{\widetilde{H}}^{p-1}\otimes\mathscr{O}(\sum m_{i}E_{i}))$.
+By induction and our choice of $p$, it is clear the morphism
+$(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p-1}\to(\pi|_{H})_{*}(\Omega_{\widetilde{H}}^{p-1}\otimes\mathscr{O}(\sum m_{i}E_{i}))$
+is surjective.
+We have also seen the morphism $\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p-1}\to(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p-1}$ is surjective for our choice of $p$.
+This produces a non-zero section of $H^{0}(\widetilde{H},\Omega_{\widetilde{H}}^{p-1}(\sum m_{i}E_{i})|_{\widetilde{H}})$ which vanishes under $\beta$.
+We can continue to iterate this process until we receive a non-zero section of $H^{0}(\widetilde{H},\mathscr{O}_{\widetilde{H}}(\sum m_{i}E_{i}))$.
+In fact, we can find for each $E_{j}$ with $m_{j}=1$ that
+$H^{0}(\widetilde{H},\mathscr{O}(m_{i}E_{i})|_{E_{j}})\neq 0.$
+This is a contradiction to the negativity lemma [GKK10, Proposition 7.5]. ∎
 This gives a different proof of Theorem 1.2 for Du Bois singularities by Theorem 1.1. We emphasize that this proof is weaker, however, since it does not imply support condition (5.1).
-6. Holomorphic Extension for (Weakly) $k$-du Bois Singularities
-6.1. Rational v.s. Weakly $k$-du Bois Singularities
+6. Holomorphic Extension for (Weakly) $m$-Du Bois Singularities
+6.1. Rational v.s. Weakly $m$-Du Bois Singularities
 A fundamental aspect of the Hodge theory of singularities is the relationship between the Du Bois property and the singularities of the MMP: Du Bois singularities are very close to log-canonical singularities, as both agree in the normal, quasi-Gorenstein case.
@@
 We have already seen a Du Bois singularity failing holomorphic extension, but they can also fail the Cohen-Macaulay property: the affine cone of a compact hyperkähler manifold in dimension $\geq 4$ is Du Bois but not Cohen-Macaulay.
-Therefore, it seems interesting to ask how closely (weakly) $k$-du Bois singularities are from having rational singularities.
+Therefore, it seems interesting to ask how closely (weakly) $m$-Du Bois singularities are from having rational singularities.
 Even for $k>0$, there is a disconnect.
@@
 Example 6.2.
-Here is an example of an affine cone singularity which is rational but is not $k$-du Bois for some $11\leq k\leq n-$ 1.
+Here is an example of an affine cone singularity which is rational but is not $m$-Du Bois for some $11\leq k\leq n-$ 1.
 The example is essentially given in [BTLM97, §4.1] and is related to the failure of Bott vanishing for non-projective spaces.
@@
 This implies that $X$ is not weakly $2$-Du Bois.
-In summary: weakly $k$-du Bois singularities need not have rational singularities, nor are rational singularities weakly $k$-du Bois for all $k$.
+In summary: weakly $m$-Du Bois singularities need not have rational singularities, nor are rational singularities weakly $m$-Du Bois for all $k$.
 What is interesting about Example 6.1 is that the 3-fold singularity is weakly $3$-Du Bois, and yet holomorphic extension does not hold in degree 3.
-This is because $\mathscr{H}^{0}\underline{\Omega}_{X}^{2}$ is also not reflexive in this case, a necessary condition for holomorphic extension to hold for $k$-du Bois singularities.
+This is because $\mathscr{H}^{0}\underline{\Omega}_{X}^{2}$ is also not reflexive in this case, a necessary condition for holomorphic extension to hold for $m$-Du Bois singularities.
 6.2. Isolated Singularities and Depth
@@
 Let $X$ be a normal variety with isolated singularities.
-If $X$ is weakly $k$-du Bois, then Theorem 2.1(vii) implies a short exact sequence
+If $X$ is weakly $m$-Du Bois, then Theorem 2.1(vii) implies a short exact sequence
 $0\to\mathscr{H}^{0}\underline{\Omega}_{X}^{p}\to\mathscr{\pi}_{*}\Omega_{\tilde{X}}^{p}\to\pi_{*}\underline{\Omega}_{E}^{p}\to 0$
@@
 The reflexivity of $\mathscr{H}^{0}\underline{\Omega}_{X}^{\dim X-1}$ appears independent of any condition on the depth, similar to the sheaf $\pi_{*}\omega_{\widetilde{X}}$.
-6.3. A criterion for Holomorphic Extension for Weakly $k$-du Bois Singularities
+6.3. A criterion for Holomorphic Extension for Weakly $m$-Du Bois Singularities
+To summarize, the weaker version of the higher Du Bois property does not appear to detect the holomorphic extension property.
+However, we can remedy this by requiring $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ to be reflexive.
+The following is an extension of Theorem 1.2 for weakly $m$-Du Bois singularities.
 Theorem 6.3.
-Suppose $X$ is a normal complex algebraic variety of dimension $n$.
-If $X$ is weakly $k$-du Bois and $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then holomorphic extension holds in degree $k+1$.
+Let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$.
+Suppose $X$ is weakly $m$-Du Bois, where $k=\mathrm{codim}_{X}(\Sigma)-1$.
+If $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then $\pi_{*}\Omega_{\widetilde{X}}^{k+1}$ is reflexive.
 Proof.
-By Proposition 1.3 and the quasi-isomorphism $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{k}$, the assumption on $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ implies $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathbb{D}_{X}(\underline{\Omega}_{X}^{k})\leq-(j+2)$ for $j\geq-n+1$.
-By looking exact triangle of Theorem 2.1(ix), dualizing, and taking cohomology, we get the vanishing $R^{n-1}\pi_{*}\Omega_{\widetilde{X}}^{n-k}(\log E)=0$.
-By the residue exact sequence, $R^{n-1}\pi_{*}\Omega_{E_{i}}^{n-k-1}=0$ for each smooth component $E_{i}$ of $E$.
-Since the problem is local, we can shrink $X$ and assume that $H^{n-1}(E_{i},\Omega_{E_{i}}^{n-k-1})=0$.
-The $E_{i}$ are smooth and projective (see §2.1), so $H^{0}(E_{i},\Omega_{E_{i}}^{k})=0$, too.
-This implies the inclusion $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{k+1})\hookrightarrow H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{k+1}(\log E))$ is an isomorphism.
-Since $X$ is du Bois, the claim follows from Theorem 1.2.
+For isolated singularities, we are checking $\pi_{*}\omega_{\widetilde{X}}$ is reflexive when $X$ is weakly $(n-1)$-Du Bois and $\mathscr{H}^{0}\underline{\Omega}_{X}^{n-1}$ is reflexive.
+By assumption, the quasi-isomorphism
+$\mathscr{H}^{0}\underline{\Omega}_{X}^{n-1}\cong\mathbf{R}\pi_{*}\Omega_{\widetilde{X}}^{n-1}(\log E)(-E)$
+implies $\dim\mathrm{Supp}~{}R^{j}\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)\leq n-j-2$.
+In particular, $R^{n-1}\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)=0$.
+This gives the additional vanishing $R^{n-1}\pi_{*}\Omega_{E_{(1)}}^{1}=0$ by the residue exact sequence.
+We may pass to cohomology by shrinking $X$ as necessary; by Hodge theory, this gives $\pi_{*}\Omega_{E_{(1)}}^{n-1}=0$.
+This implies $\pi_{*}\omega_{\widetilde{X}}\hookrightarrow\pi_{*}\omega_{\widetilde{X}}(E)$ is an isomorphism.
+Since $X$ is assumed to be Du Bois, this proves the result by Theorem 1.1.
+In order to prove $\pi_{*}\Omega_{\widetilde{X}}^{\mathrm{codim}_{X}(\Sigma)}$ is reflexive, it is sufficient to prove the stronger claim
+$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{-\dim\Sigma}^{F}\mathrm{DR}(\mathcal{IC}_{X})\leq-(j+2).$
+For $k=n-1$ (i.e., isolated singularities), this is equivalent to $\pi_{*}\omega_{\widetilde{X}}$ being reflexive.
+We can therefore proceed by induction as in the proof of Theorem 1.7.
+Specifically, the higher Du Bois property is preserved by general hyperplane [SVV23, Theorem A], and the reflexivity of $H^{0}\underline{\Omega}_{X}^{\mathrm{codim}_{X}(\Sigma)}$ is also preserved by hyperplane.
+As a result, we get the desired support condition except possibly when $j=-1$; but this is only relevant when $X$ has isolated singularities.
 ∎
 Corollary 6.4.
-Let $X$ be a normal complex variety with weakly $k$-du Bois singularities for $k\geq\mathrm{codim}_{X}(\Sigma)-1$, where $\Sigma$ is the singular locus of $X$.
+Let $X$ be a normal complex variety with weakly $m$-Du Bois singularities for $k\geq\mathrm{codim}_{X}(\Sigma)-1$, where $\Sigma$ is the singular locus of $X$.
 If $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for each $0\leq p\leq k$, then
@@
 For isolated singularities, the assumption on $k$ implies $\pi_{*}\omega_{\widetilde{X}}$ is reflexive by Theorem 6.3.
-Therefore the claim (6.1) holds by Proposition 1.3 and Grauert-Rimenschneider vanishing.
+Therefore the claim (6.1) holds by Proposition 1.3 and Grauert-Riemenschneider vanishing.
 More generally, let $X\subset Y$ be a locally closed embedding into a smooth manifold $Y$ and let $H$ be a general hyperplane section of $Y$.
-By Theorem 2.1(vi), $X\cap H$ has weakly $k$-du Bois singularities, and it is clear that $\mathscr{H}^{0}\underline{\Omega}_{X\cap H}^{p}$ is reflexive whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive.
+By Theorem 2.1(vi), $X\cap H$ has weakly $m$-Du Bois singularities, and it is clear that $\mathscr{H}^{0}\underline{\Omega}_{X\cap H}^{p}$ is reflexive whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive.
 Therefore induction and the isomorphism
@@
 6.4. A Remark on the Functorial Pullback Morphism
-Corollary 6.4 is optimal for weakly $k$-du Bois singularities.
+Corollary 6.4 is optimal for weakly $m$-Du Bois singularities.
 If $X$ is Cohen-Macaulay, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq\mathrm{codim}_{X}(\Sigma)-2$.
@@
   • If $E_{t}=\pi^{-1}(t)$ is the fiber of a resolution of singularities $\pi:\tilde{X}\to X$, then $H^{0}(E_{t},\Omega_{E_{t}}^{p}/\mathrm{tor})=0$ [Nam01, Lemma 1.2].
-Recall from §6.2 that the second condition holds for isolated weakly $k$-du Bois singularities whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq k$.
+Recall from §6.2 that the second condition holds for isolated weakly $m$-Du Bois singularities whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq k$.
 The first condition also holds in degree $p=k+1$ by Theorem 6.3.
@@
 Let $X$ be a normal complex variety with singular locus $\Sigma$, and let $\pi:\widetilde{X}\to X$ be a log-resolution of singularities.
-Suppose $X$ is weakly $k$-du Bois for $k\geq\mathrm{codim}_{X}(\Sigma)-1$ and that $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for some $p\leq k$.
-Then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p+1}$ is reflexive if and only if $\pi_{*}\Omega_{E_{i}}^{p+1}=0$ for each smooth component $E_{i}$ of the log-exceptional divisor $E$ of $\pi$.
+Suppose $X$ is weakly $m$-Du Bois for $k\geq\mathrm{codim}_{X}(\Sigma)-1$ and that $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for some $p\leq k$.
+Then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p+1}$ is reflexive if and only if the fibers $H^{0}(E_{t},\Omega_{E_{t}}^{p+1}/\mathrm{tor})=0$.
 6.5. A Remark on the Reflexivity of $\Omega_{X}^{p}$
 Let $X$ be a normal complex algebraic variety.
-We wish to consider the $k$-du Bois and $k$-rational definitions defined in [MOPW21], [JKSY21], [MP22], [FL22a], and [FL22b].
-These papers consider lci singularities, in which case the $k$-du Bois properties implies $\Omega_{X}^{p}$ is reflexive for every $p$; this is particularly special for lci singularities and restricts the codimension of the singular locus.
+We wish to consider the $m$-Du Bois and $m$-rational definitions defined in [MOPW21], [JKSY21], [MP22], [FL22a], and [FL22b].
+These papers consider lci singularities, in which case the $m$-Du Bois properties implies $\Omega_{X}^{p}$ is reflexive for every $p$; this is particularly special for lci singularities and restricts the codimension of the singular locus.
 For instance, if $Y$ is a variety with lci singularities, $\Omega_{Y}^{1}$ is reflexive if and only if $Y$ is smooth in codimension 3 [Kun86].
-Understanding what happens when $\Omega_{X}^{p}$ is reflexive seems to be a difficult problem. Instead, we consider what happens when $X$ is $k$-rational for large $k$.
-Theorem 6.6.
-If $X$ is a complex algebraic variety with gorenstein $k$-rational singularities for $k>\frac{1}{2}\dim X$, then $X$ is smooth.
+Understanding what happens when $\Omega_{X}^{p}$ is reflexive seems to be a difficult problem and rather restrictive.
+One reason for this is that the minimal generating sets of the sheaves $\Omega_{X}^{k}$ are related to the embedding dimension of $X$; in particular, a minimal generating set of the $\mathscr{O}_{X,x}$-module $\Omega_{X,x}^{k}$ has $\binom{e_{x}}{k}$ generators, where $e_{x}$ is the embedding dimension of the singularity $(X,x)$ [Gra15, §4].
+In many cases, the reflexivity of $\Omega_{X}^{k}$ necessarily restricts the embedding dimension.
+The reflexivity of the sheaf of Kähler differentials seems to be particularly special: the only known example of a singular variety with $\Omega_{X}^{p}$ reflexive for some $p\geq 1$ are locally complete intersections.
+Even in this case, the higher Kähler $p$-forms will contain torsion and cotorsion [Gra15, Theorem 1.11].
+Beyond this, little is known about the reflexivity of the sheaves of Kähler $p$-forms in general.
+We give one new result in this direction:
+Proposition 6.6.
+If $X$ is an $n$-fold Gorenstein variety with at worst quotient singularities, then $\Omega_{X}^{p}$ is reflexive for some $1\leq p\leq n$ if and only if $X$ is smooth.
 Proof.
-Let $x\in X$.
-For $p\leq k$, consider the stalk $\Omega_{X,x}^{p}$.
-Then $\mathrm{Ext}^{j}_{\mathscr{O}_{X,x}}(\Omega_{X,x}^{p},\omega_{X,x})\cong\mathscr{H}^{j}\mathbb{D}_{X}(\underline{\Omega}_{X}^{p})_{x}=0$ for $j>0$.
-This implies $\Omega_{X,x}^{p}$ is projective as an $\mathscr{O}_{X,x}$-module, and the same holds for $\mathscr{F}_{p}:=\Omega_{X,x}^{p}\otimes(\omega_{X}^{*})_{x}$ since $X$ is gorenstein.
-There is then a trace operator
-$\mathscr{F}_{p}\otimes(\mathscr{F}_{p})^{*}\to\mathscr{O}_{X,x},$
-and so any minimal generating sets of $\mathscr{F}_{p}$ and $(\mathscr{F}_{p})^{*}$ have the same cardinality, i.e., $\mathrm{rk}~{}\mathscr{F}_{p}=\mathrm{rk}~{}(\mathscr{F}_{p})^{*}$.
-Let $n=\dim X$.
-By our assumption on $k$, there is a $p<\frac{1}{2}n$ such that $\Omega_{X}^{p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ and $\Omega_{X}^{n-p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{p})$.
-Then
-$\Omega_{X}^{n-p}\cong\mathscr{H}om(\Omega_{X}^{p},\omega_{X})\cong\mathscr{H}om(\Omega_{X}^{p}\otimes\omega_{X}^{*},\mathscr{O}_{X}).$
-Using again the fact that $X$ is Gorenstein, the preceding argument says that $\mathrm{rk}~{}\Omega_{X,x}^{p}=\mathrm{rk}~{}\Omega_{X,x}^{n-p}$.
-But $\mathrm{rk}~{}\Omega_{X,x}^{r}=\binom{e_{x}}{r}$ for any $r$, where $e_{x}$ is the embedding dimension of the singularity $(X,x)$ [Gra15, §4].
-Since $e_{x}\geq n$, the equality $\binom{e_{x}}{p}=\binom{e_{x}}{n-p}$ implies that $e_{x}=n$.
-Therefore $X$ is smooth.
+We use a description of the sheaves $\pi_{*}\Omega_{\widetilde{X}}^{p}$ used in [KS21, §10].
+Let $X\subset Y$ be an embedding of $X$ into a smooth complex manifold $X$ of codimension $c$.
+Let $\mathcal{IC}_{X}$ be the $\mathscr{D}_{Y}$-module associated to the intersection Hodge module.
+The graded components of the de Rham complex with respect to the Hodge filtration are of the form
+$\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})=[\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}\xrightarrow{\nabla}\Omega_{Y}^{p+1+c}\otimes\mathrm{gr}_{c+1}^{F}\mathcal{IC}_{X}\xrightarrow{\nabla}...\xrightarrow{\nabla}\Omega_{Y}^{n+c}\otimes\mathrm{gr}_{n-p+c}^{F}\mathcal{IC}_{X}]$
+shifted by degree $-(n-p)$. Therefore, we have
+$\Omega_{X}^{[p]}\cong\mathrm{ker}(\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}\xrightarrow{\nabla}\Omega_{Y}^{p+1+c}\otimes\mathrm{gr}_{c+1}^{F}\mathcal{IC}_{X})$
+whenever $X$ has rational singularities by (3.7).
+Now we use the assumption $X$ has at worst quotient singularities. Since $X$ is a rational homology manifold, the natural morphisms
+$\underline{\Omega}_{X}^{p}\to\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})[p-n]\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$
+are quasi-isomorphisms for every $p$; in fact, there is an isomorphism
+$\mathbb{Q}_{X}[n]\xrightarrow{\sim}IC_{X}$
+of perverse sheaves, and $\mathrm{gr}_{k}^{F}\mathcal{IC}_{X}=0$ for $k\geq c+1$. The above description of $\mathrm{DR}(\mathcal{IC}_{X})$ implies
+(6.2) $\Omega_{X}^{[p]}\cong\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}.$
+If we assume $\Omega_{X}^{p}$ is reflexive, then
+$\Omega_{X}^{p}\cong\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}.$
+By Claim 10.2 and the identification (6.2), this isomorphism is closed under wedging with Kähler forms on $Y$.
+Specifically, wedging $\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}$ with Kähler $(n-p)$ forms on $Y$ lands in $\Omega_{Y}^{n+c}\otimes F_{c}\mathcal{IC}_{X}\cong\omega_{X}$.
+By restricting to $X$, we see that
+$\Omega_{X}^{n}=\Omega_{X}^{p}\wedge\Omega_{X}^{n-p}\subset\omega_{X}.$
+This implies $\Omega_{X}^{n}$ is torsion-free.
+On the other hand, since we assume $\omega_{X}$ is a line bundle, $\Omega_{X}^{n}\to\omega_{X}$ is surjective.
+Since $\omega_{X}=(\Omega_{X}^{n})^{**}$, this implies $\Omega_{X}^{n}\xrightarrow{\sim}\omega_{X}$, and so $\Omega_{X}^{n}$ is locally free.
+This means $X$ is smooth.
 ∎
-Corollary 6.7.
-If $X$ is a normal gorenstein variety with singular locus $\Sigma$. If $X$ is $k$-rational with $k>\frac{1}{2}\mathrm{codim}_{X}(\Sigma)$, then $X$ is smooth.
-Proof.
-The case of isolated singularities is covered by Theorem 6.6. The more general fact follows since a general hyperplane section of a $k$-rational singularity is $k$-rational [SVV23, Theorem A]. ∎
-We remark that Corollary 6.7 was observed for the $k$-du Bois property in the case of hypersurface singularities [MOPW21, Lemma 2.2], which is equivalent to being $(k-1)$-rational [FL22b, Corollary 1.9].
-We can recover this result for isolated singularities by the Proof of Theorem 6.6, since we only require:
-  • $X$ is $k$-rational for some $k\leq\frac{1}{2}\dim X$,
-  • $\Omega_{X}^{n-k}$ is reflexive.
-While $\Omega_{X}^{n-k}$ may not be reflexive if $X$ is only $k$-du Bois, this always holds for hypersurface singularities.
-7. Hodge Theory of Proper Varieties with Weakly $k$-Rational Singularities
+7. Hodge Theory of Proper Varieties with Weakly $m$-rational Singularities
 7.1. Mixed Hodge Theory for MMP Singularities
@@
 In his thesis, Deligne defined the notion of a mixed Hodge structure and constructed a canonical mixed Hodge structure on the cohomology of any complex algebraic variety [Del71], [Del74].
-Specfically, there is an increasing weight filtration $W^{\bullet}$ on $H^{k}(X,\mathbb{Q})$ and a decreasing filtration $F^{\bullet}$ on $H^{k}(X,\mathbb{C})$ which descends to a Hodge filtration on $\mathrm{gr}_{p}^{W}H^{k}(X,\mathbb{Q})\otimes\mathbb{C}$ for each $k$.
+Specfically, there is an increasing weight filtration $W_{\bullet}$ on $H^{k}(X,\mathbb{Q})$ and a decreasing filtration $F^{\bullet}$ on $H^{k}(X,\mathbb{C})$ which descends to a Hodge filtration on $\mathrm{gr}_{p}^{W}H^{k}(X,\mathbb{Q})\otimes\mathbb{C}$ for each $k$.
 The Du Bois complex is a byproduct of this construction: namely, Du Bois details in [DB81] that Deligne’s construction produces a complex $\underline{\Omega}_{X}^{\bullet}$ for any algebraic variety $X$ which generates the Hodge filtration under the spectral sequence
@@
 For $k=2$, this map is again injective by Leray when $X$ has rational singularities.
-The obvious question to ask is how does the (weakly) $k$-du Bois property affect the mixed Hodge theory on the cohomology $H^{*}(X,\mathbb{Q})$ of a projective variety.
+The obvious question to ask is how does the (weakly) $m$-Du Bois property affect the mixed Hodge theory on the cohomology $H^{*}(X,\mathbb{Q})$ of a projective variety.
 Surprisingly, the vanishing cohomology of the $\underline{\Omega}_{X}^{p}$ has little control over the weight filtration:
 Example 7.1.
-  1. (1)
+  1. 7.1.
     Let $X$ be a projective curve.
@@
 For an explicit example, consider the nodal elliptic curve.
-  2. (2)
+  2. 7.2.
     Let $X$ be a projective hypersurface of dimension 3 with ordinary double points.
@@
 Therefore, $H^{2}(X,\mathbb{Q})$ carries a pure Hodge structure by the preceding discussion, but $H^{3}(X,\mathbb{Q})$ need not carry a pure Hodge structure
-  3. (3)
+  3. 7.3.
     Let $Y$ be a projective hyperkähler 4-fold manifold.
 Suppose there exists a birational contraction $\phi:Y\to X$ of a Lagrangian submanifold $L\cong\mathbb{P}^{2}\subset Y$ to a point.
-Then $X$ is not weakly $k$-du Bois for $k>1$, but $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure for each $k$.
+Then $X$ is not weakly $1$-rational [Tig24, Proposition 1.4], but $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure for each $k$.
 Moreover, $H^{k}(X,\mathbb{Q})$ will carry a pure Hodge structure for large $k$ for topological reasons: if $X$ is projective and has isolated singularities, then $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure of weight $k$ for $k>n$.
@@
 More generally, we can consider the weight filtrations for the mixed Hodge structures on $H^{k}(X,\mathbb{Q})$ and $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$, respectively.
-Since $X$ is proper, the weight filtration $W^{\bullet}$ truncates to $H^{k}(X,\mathbb{Q})$ for each $k$.
+Since $X$ is proper, the weight filtration $W_{\bullet}$ truncates to $H^{k}(X,\mathbb{Q})$ for each $k$.
 The weight filtration on the dual $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$ therefore is supported in higher weights (compare this to the cohomology $H^{k}(X_{\mathrm{reg}},\mathbb{Q})$), and these groups are related to the Grothendieck duals $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ of the Du Bois complex.
-It therefore seems better to consider how the $k$-rational property affects the weight filtration.
-7.2. Purity for $k$-Rational Singularities
+It therefore seems better to consider how the $m$-rational property affects the weight filtration.
+7.2. Purity for $m$-rational Singularities
 Theorem 7.2.
-If $X$ is a projective variety of dimension $n$ with weakly $k$-rational singularities, then the canonical mixed Hodge structure on $H^{k}(X,\mathbb{Q})$ is pure of weight $k$.
+If $X$ is a normal and projective variety of dimension $n$ with weakly $m$-rational singularities, then the canonical mixed Hodge structure on $H^{m}(X,\mathbb{Q})$ is pure of weight $m$.
 Proof.
@@
 $\Omega_{X}^{[p]}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$
-for every $0\leq p\leq k$.
+for every $0\leq p\leq m$.
 On the one hand, the spectral sequence
 $E_{1}^{p,q}=\mathbb{H}^{q}(X,\underline{\Omega}_{X}^{p})\Rightarrow H^{p+q}(X,\mathbb{C})$
-generates the Hodge filtration. Note that the weight filtration $W_{\mathbb{C}}^{\bullet}$ is supported in lower weights, since $X$ is proper.
+generates the Hodge filtration. Note that the weight filtration $W_{\mathbb{C}}^{\bullet}$ is supported in weight $\leq n$, since $X$ is proper.
 On the other hand, $\mathbb{H}^{q}(X,\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p}))\cong\mathrm{Hom}_{\mathscr{O}_{\mathrm{pt}}}(\mathbb{H}^{n-q}(X,\underline{\Omega}_{X}^{n-p}),\mathscr{O}_{\mathrm{pt}})$ by duality.
-Therefore $E_{1}^{p,q}$ generates the Hodge filtration on the dual mixed Hodge structure $H^{n-k}(X,\mathbb{C})^{*}$ for $k=p+q$.
-The weight filtration on this mixed Hodge structure is supported in higher weights.
-Therefore $H^{k}(X,\mathbb{Q})$ must be a pure Hodge structure.
+Therefore $E_{1}^{p,q}$ generates the Hodge filtration on the dual mixed Hodge structure $H^{n-m}(X,\mathbb{C})^{*}$ for $m=p+q$.
+The weight filtration on this mixed Hodge structure is supported in weight $\geq n$.
+Therefore, $H^{m}(X,\mathbb{Q})$ must be a pure Hodge structure.
 ∎
 Corollary 7.3.
-If $X$ is a projective variety of dimension $n$ with weakly $k$-rational singularities, then there is a non-canonical decomposition
-$H^{k}(X,\mathbb{C})=\bigoplus_{p+q=k}H^{q}(X,\pi_{*}\Omega_{\tilde{X}}^{p})$
+If $X$ is a projective variety of dimension $n$ with weakly $m$-rational singularities, then there is a non-canonical decomposition
+$H^{m}(X,\mathbb{C})=\bigoplus_{p+q=m}H^{q}(X,\pi_{*}\Omega_{\tilde{X}}^{p})$
 induced by the Hodge filtration.
-We remark that Theorem 7.2 implies something stronger than the cohomology $H^{k}(X,\mathbb{Q})$ carrying a pure Hodge structure.
-For instance if $k<\mathrm{codim}_{X}(\Sigma)$, where $\Sigma$ is the singular locus, the above proof shows that $H^{k}(X,\mathbb{Q})\to H^{k}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism.
-If $X$ has rational singularities, this morphism is always injective for $k=2$.
-If $X$ is a 3-fold with isolated singularities, then $H^{2}(X,\mathbb{Q})\to H^{2}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism if and only if Poincaré duality holds: therefore the defect $\sigma(X)$ of $X$ must be 0 [Kaw88, p.
+We remark that Theorem 7.2 implies something stronger than the cohomology $H^{m}(X,\mathbb{Q})$ carrying a pure Hodge structure.
+For instance if $m<\mathrm{codim}_{X}(\Sigma)$, where $\Sigma$ is the singular locus, the above proof shows that $H^{m}(X,\mathbb{Q})\to H^{m}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism.
+If $X$ has rational singularities, this morphism is always injective for $m=2$.
+If $X$ is a 3-fold with isolated singularities, then $H^{m}(X,\mathbb{Q})\to H^{m}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism if and only if Poincaré duality holds: therefore the defect $\sigma(X)$ of $X$ must be 0 [Kaw88, p.
 97], and $X$ is $\mathbb{Q}$-factorial.